### v4

# Universidade de Lisboa, Av. Rovisco Pais 1, 1049-001 Lisboa, Portugal

Abstract
We introduce a two real scalar singlet extension of the two Higgs doublet model. We study the vacuum structure, the bounded from below conditions, the restrictions from the oblique parameters S,T and U, as well as the unitarity constraints. We submit the model to collider and Dark Matter experimental constraints and explore its allowed parameter space. We compare randomly populated simulations, simulations starting near the alignment limit, and a Machine Learning based exploration. Using Evolutionary Strategies, we efficiently search for regions with a viable Dark Matter candidate.
1 Introduction
Despite the large success of the Standard Model (SM) of particle physics in providing experimental predictions, leading to the discovery of a new scalar particle resembling the predicted Higgs boson Aad and others (2012); Chatrchyan and others (2012), there is a general consensus that there must be Physics Beyond the Standard Model (BSM). The critical open problems consist of: the need for new sources of CP-violation, a necessary ingredient for a successful explanation of the baryon asymmetry of the universe Sakharov (1967); an explanation for the existence and nature of Dark Matter (DM), which comprises of order 85 % of the matter content of the Universe Ade and others (2014); and the origin of the observed tiny neutrino masses. Motivated by the lack of a fundamental reason why the scalar sector should be limited to a single Higgs doublet, Higgs-sector extensions are required for many of the viable explanations to these problems.
In this work, we consider a two Higgs doublet model (2HDM) Branco et al. (2012) with a $\mathbbm{Z}_{2}$ symmetry imposed in order to forbid Higgs-mediated flavor changing neutral couplings (FCNCs) at the tree-level, with fermions of a given electric charge coupling to only one Higgs doublet. Of the four possible choices, the type II 2HDM is the most studied, since it corresponds to the structure present in supersymmetric models - see, for example, refs. Djouadi (2008); Dreiner et al. (2023). We, however, consider the general type II 2HDM and aim to extend the scalar sector with additional singlets as viable candidates for particle Dark Matter. When protected by a new global symmetry that remains unbroken in the vacuum, the singlets can meet the requirements of heavy, stable, electrically neutral particles of non-baryonic nature Dimopoulos et al. (1990); McDermott et al. (2011). We will consider a scenario where the DM particle are produced via the freeze-out mechanism Bertone et al. (2005); Feng (2010). A particle candidate with a mass similar to the scale of electroweak symmetry breaking and an interaction cross section with the SM particles of the order of the weak force processes can meet the observed relic abundance, and fits the class of Weakly Interacting Massive Particles (WIMPs). The model is then subjected to all theoretical, collider, astrophysical, and cosmological constraints to obtain the allowed parameter space region.
The 2HDM extended by an additional real singlet with a $\mathbbm{Z}_{2}$ parity symmetry may contain a viable Dark Matter candidate when the singlet does not acquire a vacuum expectation value (vev), see e.g. refs. Grzadkowski and Osland (2010); Logan (2011); Boucenna and Profumo (2011); He et al. (2012); Bai et al. (2013); Drozd et al. (2014); Campbell et al. (2015). It has been dubbed the Next-to-Two Higgs Doublet model (N2HDM) in Chen et al. (2014); Muhlleitner et al. (2017); Ferreira et al. (2019); Engeln et al. (2020); Glaus et al. (2022). In 2022, there was already a tension in the N2HDM between the relic density measured by PLANCK Aghanim and others (2020) and exclusion bounds from DM direct detection experiments, as shown in figure 3 of Glaus et al. (2022). Since then, the situation has worsened, given the great improvement in the scattering bounds Aalbers and others (2024a).
A possible alternative is to consider a Type II 2HDM augmented with a complex scalar singlet (2HDMS), either without a DM candidate Baum and Shah (2018); Heinemeyer et al. (2022), or with a DM candidate enabled through a suitable symmetry and vacuum Dutta et al. (2025, 2024); Lahiri and Moortgat-Pick (2024). Such studies have concentrated thus far on benchmarks rather than dedicated multi-variable scans.
Extensions of the scalar sector of the 2HDM through the addition of more than one singlet offer promising frameworks that could also address the Dark Matter puzzle  [^1] The interesting possibility of adding multiple singlets to the (one Higgs) SM, has also been considered. For a recent analysis see, for example, Ref. Gonçalves et al. (2025).. Within this context, we consider a two real scalar inert singlet extension of the two Higgs doublet model (2HDM2S). It includes all theoretical derivations relevant for unitarity, boundedness from below, and vacuum stability. It also includes a full parameter scan simulation of the model, performed in three ways: i) with a traditional scan of the full parameter space; ii) starting from scans close to the alignment limit; and, iii) using a Machine Learning optimization technique, evoking Evolutionary Strategies.
The work’s structure is as follows. In section 2, the scalar potential of the model and its particle content are developed. In section 3, the vacuum structure and the interplay of multiple vacua is analysed. The boundedness from below sufficient conditions, the expressions for the potential at the minima, the perturbative unitarity constraints, and EW precision observables constraints for this model are developed in sections 4, 5, 6, and 7, respectively. An overview of the experimental constraints applied to this model is given in section 8. In section 9, the computational methods used are explained, and a comparison of results between different approaches is done. A discussion of results satisfying all described constraints is provided in section 10, followed by conclusions in section 11. The appendix shows a comparison between the desired vacuum and all other possible vacua within this model.
2 The Scalar Potential of the 2HDM2S
Our goal is to analyze an extension to the SM, with two doublets and two real scalars, in which the scalars do not acquire a Vacuum Expectation Value (vev) after Spontaneous Symmetry Breaking (SSB). In order to avoid dangerous flavor-changing neutral couplings (FCNC) at the tree level, a $\mathbbm{Z}_{2}$ symmetry softly broken by the $m_{12}^{2}$ term, to introduce a decoupling limit Gunion and Haber (2003), is imposed on the Lagrangian. In addition, we assume CP conservation in the theory; therefore, all the coefficients in the potential are taken to be real.
The 2HDM2S contains an additional complex doublet to the Standard Model, $\Phi_{2}$, and two real scalars, S and P, giving rise to additional terms in the scalar potential, including one $\mathbb{Z}^{\prime}_{2}$ symmetry imposed on the real scalar singlet fields S and P. The imposed symmetries are
$\mathbb{Z}_{2}:\;\;\;\;\Phi_{1}\rightarrow\Phi_{1},\;\;\;\;\Phi_{2}\rightarrow-\Phi_{2},\;\;\;\;S\rightarrow S,\;\;\;\;P\rightarrow P,$ (1)
and
$\mathbb{Z}_{2}^{\prime}:\;\;\;\;\Phi_{1}\rightarrow\Phi_{1},\;\;\;\;\Phi_{2}\rightarrow\Phi_{2},\;\;\;\;S\rightarrow-S,\;\;\;\;P\rightarrow-P.$ (2)
The terms invariant under the symmetry are written as,
$V=$ $\,m_{11}^{2}(\Phi_{1}^{\dagger}\Phi_{1})+m_{22}^{2}(\Phi_{2}^{\dagger}\Phi_{2})-\big[m_{12}^{2}(\Phi_{1}^{\dagger}\Phi_{2})+\text{h.c.}\big]+\frac{\lambda_{1}}{2}(\Phi_{1}^{\dagger}\Phi_{1})^{2}+\frac{\lambda_{2}}{2}(\Phi_{2}^{\dagger}\Phi_{2})^{2}$ (3)
$+\lambda_{3}(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{2}^{\dagger}\Phi_{2})+\lambda_{4}(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{2}^{\dagger}\Phi_{1})+\bigg[\frac{\lambda_{5}}{2}(\Phi_{1}^{\dagger}\Phi_{2})^{2}+\text{h.c.}\bigg]$
$\,\frac{1}{2}m_{S}^{2}S^{2}+\frac{1}{2}m_{P}^{2}P^{2}-m_{SP}^{2}SP+\frac{1}{8}\lambda_{6}S^{4}+\frac{1}{8}\lambda_{9}P^{4}+\frac{1}{4}\lambda_{10}S^{2}P^{2}+\frac{1}{6}\lambda_{13}S^{3}P$
$+\frac{1}{6}\lambda_{14}SP^{3}+\frac{1}{2}\big(\lambda_{7}(\Phi_{1}^{\dagger}\Phi_{1})+\lambda_{8}(\Phi_{2}^{\dagger}\Phi_{2})\big)S^{2}+\frac{1}{2}\big(\lambda_{11}(\Phi_{1}^{\dagger}\Phi_{1})+\lambda_{12}(\Phi_{2}^{\dagger}\Phi_{2})\big)P^{2}$
$+\frac{1}{2}\big(\lambda_{15}(\Phi_{1}^{\dagger}\Phi_{1})+\lambda_{16}(\Phi_{2}^{\dagger}\Phi_{2})\big)SP.$
We can alternatively write potential of the 2HDM2S as,
$V=V_{2}+V_{4}\ ,$ (4)
where the quadratic part is
$V_{2}=m_{11}^{2}(\Phi_{1}^{\dagger}\Phi_{1})+m_{22}^{2}(\Phi_{2}^{\dagger}\Phi_{2})-\left(m_{12}^{2}(\Phi_{1}^{\dagger}\Phi_{2})+\text{h.c.}\right)+\frac{1}{2}m_{S}^{2}S^{2}+\frac{1}{2}m_{P}^{2}P^{2}-m_{SP}^{2}SP\,.$ (5)
and the quartic terms are,
$V_{\rm Quartic}=V_{N}+V_{CB}+V_{HC}+V_{O}\,$, (6)
where
$V_{N}=$ $\frac{\lambda_{1}}{2}(\Phi_{1}^{\dagger}\Phi_{1})^{2}+\frac{\lambda_{2}}{2}(\Phi_{2}^{\dagger}\Phi_{2})^{2}+(\lambda_{3}+\lambda_{4})(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{2}^{\dagger}\Phi_{2})+\frac{1}{8}\lambda_{6}S^{4}+\frac{1}{8}\lambda_{9}P^{4}$
$+\frac{1}{4}\lambda_{10}S^{2}P^{2}+\frac{1}{2}\left(\lambda_{7}(\Phi_{1}^{\dagger}\Phi_{1})+\lambda_{8}(\Phi_{2}^{\dagger}\Phi_{2})\right)S^{2}+\frac{1}{2}\left(\lambda_{11}(\Phi_{1}^{\dagger}\Phi_{1})+\lambda_{12}(\Phi_{2}^{\dagger}\Phi_{2})\right)P^{2}\,$, (7)
$V_{CB}=$ $-\lambda_{4}z_{12}\,$, (8)
$V_{HC}=$ $\frac{1}{2}\lambda_{5}\left((\Phi_{1}^{\dagger}\Phi_{2})^{2}+\text{h.c.}\right)\,$, (9)
$V_{O}=$ $\frac{1}{6}\lambda_{13}S^{3}P+\frac{1}{6}\lambda_{14}SP^{3}+\frac{1}{2}\left(\lambda_{15}(\Phi_{1}^{\dagger}\Phi_{1})+\lambda_{16}(\Phi_{2}^{\dagger}\Phi_{2})\right)SP\,$, (10)
where we have defined Faro and Ivanov (2019),
$z_{12}=(\Phi_{1}^{\dagger}\Phi_{1})(\Phi_{2}^{\dagger}\Phi_{2})-(\Phi_{1}^{\dagger}\Phi_{2})(\Phi_{2}^{\dagger}\Phi_{1}).$ (11)
We consider the vacuum where the singlets do not develop a vev, thus keeping the $\mathbb{Z}_{2}^{\prime}$ symmetry intact. After SSB, we can develop the doublets and singlets in the symmetry basis as:
$\Phi_{1}=\begin{pmatrix}w_{1}^{+}\\ \frac{(v_{1}+y_{1}+iz_{1})}{\sqrt{2}}\end{pmatrix},\;\;\;\;\Phi_{2}=\begin{pmatrix}w_{2}^{+}\\ \frac{(v_{2}+y_{2}+iz_{2})}{\sqrt{2}}\end{pmatrix},\;\;\;\;S=s,\;\;\;\;P=p.$ (12)
The particle content of the model consists of four CP-even Higgs bosons, $h_{i}$, $i\in\{1,2,3,4\}$, where $h_{3}$ and $h_{4}$ are on the dark sector, a pseudoscalar Higgs, $A$, and a charged Higgs $H^{\pm}$.
Due to charge and CP conservation and the imposed symmetries, the $8\times 8$ mass matrix is decomposed into four blocks: the $2\times 2$ matrix for the charged fields, the $2\times 2$ matrix for the CP-odd fields and two $2\times 2$ matrices for the CP-even states, since the fields from the doublets do not mix with the scalars $s$ and $p$. The charged and pseudoscalar sectors remain unchanged in regard to the 2HDM. Hence, the charged and pseudoscalar mass matrices may be diagonalized by the rotation matrix
$\mathcal{O}_{\beta}=\begin{pmatrix}c_{\beta}&s_{\beta}\\ -s_{\beta}&c_{\beta}\end{pmatrix}$, (13)
where $\tan\beta=\frac{v_{2}}{v_{1}}$, $s_{\beta}=\sin(\beta)$, and $c_{\beta}=\cos(\beta)$. With this definition we may obtain the physical-mass eigenstates Das and Saha (2019) by
$\begin{bmatrix}G^{0}\\ A\end{bmatrix}=\mathcal{O}_{\beta}\begin{bmatrix}z_{1}\\ z_{2}\end{bmatrix},\quad\begin{bmatrix}G^{+}\\ H^{+}\end{bmatrix}=\mathcal{O}_{\beta}\begin{bmatrix}w_{1}^{+}\\ w_{2}^{+}\end{bmatrix}$, (14)
where $G^{+}$ and $G^{0}$ are the Nambu-Goldstone bosons. For the scalar states, we obtain
$\begin{bmatrix}h_{1}\\ h_{2}\end{bmatrix}=\begin{bmatrix}c_{\alpha_{1}}&-s_{\alpha_{1}}\\ s_{\alpha_{1}}&c_{\alpha_{1}}\end{bmatrix}\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix},\\ \begin{bmatrix}h_{3}\\ h_{4}\end{bmatrix}=\begin{bmatrix}c_{\alpha_{2}}&-s_{\alpha_{2}}\\ s_{\alpha_{2}}&c_{\alpha_{2}}\end{bmatrix}\begin{bmatrix}s\\ p\end{bmatrix}$. (15)
Fixing the mass basis, one derives the corresponding 22 free parameters:
$m_{h1,h2,h3,h4},\;m_{A},\;m_{H\pm},\;m_{12}^{2},\;m_{S}^{2},\;m_{P}^{2},\;m_{SP}^{2},\;\lambda_{6,8,9,10,12,13,14,16},\;\alpha_{1,2},\;t_{\beta},\;v\;.$ (16)
We may now write the quartic couplings $\lambda_{i}$ in terms of the physical basis, obtaining the following expressions:
$\tilde{\mu}^{2}=$ $\frac{m_{12}^{2}}{s_{\beta}c_{\beta}},$ (17)
$\lambda_{1}=$ $\frac{1}{v^{2}c_{\beta}^{2}}\big(m_{1}^{2}c_{\alpha_{1}}^{2}+m_{2}^{2}s_{\alpha_{1}}^{2}-\tilde{\mu}^{2}s_{\beta}^{2}\big),$ (18)
$\lambda_{2}=$ $\frac{1}{v^{2}s_{\beta}^{2}}\big(m_{1}^{2}s_{\alpha_{1}}^{2}+m_{2}^{2}c_{\alpha_{1}}^{2}-\tilde{\mu}^{2}c_{\beta}^{2}\big),$ (19)
$\lambda_{3}=$ $\frac{1}{v^{2}}\big(\frac{s_{\alpha_{1}}c_{\alpha_{2}}}{s_{\beta}c_{\beta}}\big(m_{2}^{2}-m_{1}^{2}\big)-\tilde{\mu}^{2}+2m^{2}_{H\pm}\big),$ (20)
$\lambda_{4}=$ $\frac{1}{v^{2}}\big(\tilde{\mu}^{2}+m_{A}^{2}-2m^{2}_{H\pm}\big),$ (21)
$\lambda_{5}=$ $\frac{1}{v^{2}}\big(\tilde{\mu}^{2}-m_{A}^{2}\big),$ (22)
$\lambda_{7}=$ $\frac{1}{v^{2}c_{\beta}^{2}}\big(2m_{DMS}^{2}c_{\alpha_{2}}^{2}+2m_{DMP}^{2}s_{\alpha_{2}}^{2}-2m_{S}^{2}-\lambda_{8}v^{2}s_{\beta}^{2}\big),$ (23)
$\lambda_{11}=$ $\frac{1}{v^{2}c_{\beta}^{2}}\big(2m_{DMS}^{2}s_{\alpha_{2}}^{2}+2m_{DMP}^{2}c_{\alpha_{2}}^{2}-2m_{P}^{2}-\lambda_{12}v^{2}s_{\beta}^{2}\big),$ (24)
$\lambda_{15}=$ $\frac{1}{v^{2}c_{\beta}^{2}}\big[4\big(m_{DMP}^{2}-m_{DMS}^{2}\big)s_{\alpha_{2}}c_{\alpha_{2}}+4m_{SP}^{2}-\lambda_{16}v^{2}s_{\beta}^{2}\big].$ (25)
3 Vacuum Structure
3.1 Stationarity Conditions
The most general constant field configuration for the vacuum is, using the notation from Muhlleitner et al. (2017); Ferreira et al. (2019),
$\langle\Phi_{1}\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v_{1}\end{pmatrix},\;\;\;\;\langle\Phi_{2}\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}v_{cb}\\ v_{2}+iv_{cp}\end{pmatrix},\;\;\;\;\langle S\rangle=v_{S},\;\;\;\;\langle P\rangle=v_{P}.$ (26)
In order to find all the possible minima, we consider the following stationarity conditions for the vevs or, equivalently, for the complex charged fields $w_{i}^{+}\;(i\in\{1,2\})$, the real neutral CP-even, $y_{1},y_{2},s$, and $p$, and CP-odd fields, $z_{1}$, and $z_{2}$ :
$\left\langle\frac{\partial V}{\partial w_{1}^{+}}\right\rangle=0\iff$ $v_{cb}\big(v_{1}v_{2}\big(\lambda_{4}+\lambda_{5}\big)-2m_{12}^{2}\big)=0,$ (27)
$\left\langle\frac{\partial V}{\partial w_{2}^{+}}\right\rangle=\left\langle\frac{\partial V}{\partial v_{cb}}\right\rangle=0\iff$ $-v_{cb}m_{12}^{2}=\frac{1}{2}v_{cb}\big(v_{1}^{2}\lambda_{3}+v_{2}^{2}\lambda_{3}+v_{cb}^{2}\lambda_{2}+v_{cp}^{2}\lambda_{2}+v_{S}^{2}\lambda_{8}$
$+v_{P}^{2}\lambda_{12}+v_{P}v_{S}\lambda_{16}\big),$ (28)
$\left\langle\frac{\partial V}{\partial y_{1}}\right\rangle=\left\langle\frac{\partial V}{\partial v_{1}}\right\rangle=0\iff$ $v_{2}m_{12}^{2}-v_{1}m_{1}^{2}=\frac{1}{2}v_{1}\big(v_{1}^{2}\lambda_{1}+v_{2}^{2}\lambda_{345}+v_{cb}^{2}\lambda_{3}+v_{cp}^{2}\lambda_{34-5}$
$+v_{S}^{2}\lambda_{7}+v_{P}^{2}\lambda_{11}+v_{S}v_{P}\lambda_{15}\big),$ (29)
$\left\langle\frac{\partial V}{\partial y_{2}}\right\rangle=\left\langle\frac{\partial V}{\partial v_{2}}\right\rangle=0\iff$ $v_{1}m_{12}^{2}-v_{2}m_{2}^{2}=\frac{1}{2}v_{2}\big(v_{1}^{2}\lambda_{345}+v_{2}^{2}\lambda_{2}+v_{cb}^{2}\lambda_{2}+v_{cp}^{2}\lambda_{2}$
$+v_{S}^{2}\lambda_{8}+v_{P}^{2}\lambda_{12}+v_{S}v_{P}\lambda_{16}\big),$ (30)
$\left\langle\frac{\partial V}{\partial z_{1}}\right\rangle=0\iff$ $v_{cp}\big(v_{1}v_{2}\lambda_{5}-2m_{12}^{2}\big)=0,$ (31)
$\left\langle\frac{\partial V}{\partial z_{2}}\right\rangle=\left\langle\frac{\partial V}{\partial v_{cp}}\right\rangle=0\iff$ $-v_{cp}m_{12}^{2}=\frac{1}{2}v_{cp}\big(v_{1}^{2}\lambda_{34-5}+v_{2}^{2}\lambda_{2}+v_{cb}^{2}\lambda_{2}+v_{cp}^{2}\lambda_{2}$
$+v_{S}^{2}\lambda_{8}+v_{P}^{2}\lambda_{12}+v_{S}v_{P}\lambda_{16}\big),$ (32)
$\left\langle\frac{\partial V}{\partial s}\right\rangle=\left\langle\frac{\partial V}{\partial v_{S}}\right\rangle=0\iff$ $v_{P}m_{SP}^{2}-v_{S}m_{S}^{2}=\frac{1}{2}v_{S}\big(v_{1}^{2}\lambda_{7}+v_{2}^{2}\lambda_{8}+v_{cb}^{2}\lambda_{8}+v_{cp}^{2}\lambda_{8}$
$+v_{S}^{2}\lambda_{6}+v_{P}^{2}\lambda_{10}+v_{S}v_{P}\lambda_{13}\big)+\frac{1}{4}v_{P}\big(v_{1}^{2}\lambda_{15}+v_{2}^{2}\lambda_{16}$
$+v_{cb}^{2}\lambda_{16}+v_{cp}^{2}\lambda_{16}\big)+\frac{1}{6}v_{P}^{3}\lambda_{14},$ (33)
$\left\langle\frac{\partial V}{\partial p}\right\rangle=\left\langle\frac{\partial V}{\partial v_{P}}\right\rangle=0\iff$ $v_{S}m_{SP}^{2}-v_{P}m_{P}^{2}=\frac{1}{2}v_{P}\big(v_{1}^{2}\lambda_{11}+v_{2}^{2}\lambda_{12}+v_{cb}^{2}\lambda_{12}+v_{cp}^{2}\lambda_{12}$
$+v_{S}^{2}\lambda_{10}+v_{P}^{2}\lambda_{9}+v_{S}v_{P}\lambda_{14}\big)+\frac{1}{4}v_{S}\big(v_{1}^{2}\lambda_{15}+v_{2}^{2}\lambda_{16}$
$+v_{cb}^{2}\lambda_{16}+v_{cp}^{2}\lambda_{16}\big)+\frac{1}{6}v_{S}^{3}\lambda_{13},$ (34)
where
$\lambda_{345}\equiv\lambda_{3}+\lambda_{4}+\lambda_{5},\;\;\;\;\lambda_{34-5}\equiv\lambda_{3}+\lambda_{4}-\lambda_{5}.$ (35)
From eqs. (28) and (32) we conclude that
$\big(\lambda_{4}=\lambda_{5})\vee\big[\big(v_{cb}\neq 0\Rightarrow v_{cp}=0\big)\vee\big(v_{cp}\neq 0\Rightarrow v_{cb}=0\big)\big].$ (36)
From eqs. (29) and (30) we observe that
$\big(v_{1}=0\Leftrightarrow v_{2}=0\big)\vee m_{12}^{2}=0.$ (37)
From eqs. (27) and (31), we further infer that
$v_{1}=v_{2}=0\Rightarrow\big(v_{cb}=v_{cp}=0\vee m_{12}^{2}=0\big).$ (38)
These results are similar to those on N2HDM.
From the eigenvalues of the scalar mass matrices, we derive, at a given stationary point $i$, the squared charged scalar mass and the pseudoscalar squared mass (for $v_{cb}=0\wedge v_{vp}=0$)
$\big(m_{H\pm}^{2}\big)_{i}=m_{12}^{2}\frac{v_{i}^{2}}{(v_{1})_{i}(v_{2})_{i}}-\frac{1}{2}(\lambda_{4}+\lambda_{5})v_{i}^{2},\;\;\;\big(m_{A}^{2}\big)_{i}=\big(m_{H\pm}^{2}\big)_{i}+\frac{1}{2}(\lambda_{4}-\lambda_{5})v_{i}^{2}.$ (39)
From the scalar mass matrix of the singlets, we define the squared mass of the singlet field S and P without mass mixing as $\big(m_{s}^{2}\big)_{i}$ and $\big(m_{p}^{2}\big)_{i}$ which, due to the required copositivity of the CP-even mass matrix, are positive. For the mass mixing term (off-diagonal), we define it as $\big(m_{sp}^{2}\big)_{i}$, which is not necessarily positive.
3.2 Formalism and Vacua
To study the interplay of multiple vacua, it is useful to introduce a bilinear formalism, similar to the one used in Lahiri and Moortgat-Pick (2024); Ferreira et al. (2019) for the 2HDMS and N2HDM. The relevant bilinears for this work are the following.
$x_{1}=|\Phi_{1}|^{2},x_{2}=|\Phi_{2}|^{2},x_{3}=\operatorname{Re}\big(\Phi_{1}^{\dagger}\Phi_{2}\big),x_{4}=\operatorname{Im}\big(\Phi_{1}^{\dagger}\Phi_{2}\big),x_{5}=\frac{1}{2}S^{2},x_{6}=\frac{1}{2}P^{2},x_{7}=\frac{1}{2}SP.$ (40)
We define the vectors X, A and the symmetric matrix B as
$X=\begin{pmatrix}x_{1}\\ x_{2}\\ x_{3}\\ x_{4}\\ x_{5}\\ x_{6}\\ x_{7}\end{pmatrix},\;\;A=\begin{pmatrix}m_{11}^{2}\\ m_{22}^{2}\\ -2m_{12}^{2}\\ 0\\ m_{S}^{2}\\ m_{P}^{2}\\ -2m_{SP}^{2}\end{pmatrix},\;\;B=\begin{pmatrix}\lambda_{1}&\lambda_{3}&0&0&\lambda_{7}&\lambda_{11}&\lambda_{15}\\ \lambda_{3}&\lambda_{2}&0&0&\lambda_{8}&\lambda_{12}&\lambda_{16}\\ 0&0&2(\lambda_{4}+\lambda_{5})&0&0&0&0\\ 0&0&0&2(\lambda_{4}-\lambda_{5})&0&0&0\\ \lambda_{7}&\lambda_{8}&0&0&\lambda_{6}&\lambda_{10}&\frac{2}{3}\lambda_{13}\\ \lambda_{11}&\lambda_{12}&0&0&\lambda_{10}&\lambda_{9}&\frac{2}{3}\lambda_{14}\\ \lambda_{15}&\lambda_{16}&0&0&\frac{2}{3}\lambda_{13}&\frac{2}{3}\lambda_{14}&0\end{pmatrix}$. (41)
In terms of the bilinears, the potential may be written as
$V=A^{T}X+\frac{1}{2}X^{T}BX.$ (42)
We also make use of the vector
$V^{\prime}=\frac{\partial V}{\partial X^{T}}=A+BX,$ (43)
so that, at a stationary point $i$, the value of the potential is given by
$V_{i}=\frac{1}{2}A^{T}\langle X\rangle_{i}=-\frac{1}{2}\langle X\rangle_{i}^{T}B\langle X\rangle_{i}.$ (44)
The procedure for finding the stability conditions between two stationary points $i$ and $j$ is as follows. The internal product between $X_{i}$ and $V^{\prime}_{j}$ yields
$\langle X\rangle_{i}^{T}V^{\prime}_{j}=\langle X\rangle_{i}^{T}A+\langle X\rangle_{i}^{T}B\langle X\rangle_{j}\,$, (45)
and the internal product between $X_{j}$ and $V^{\prime}_{i}$ results in
$\langle X\rangle_{j}^{T}V^{\prime}_{i}=\langle X\rangle_{j}^{T}A+\langle X\rangle_{j}^{T}B\langle X\rangle_{i}.$ (46)
From eq. (44), one can write
$\langle X\rangle_{i}^{T}A=2V_{i},\;\;\;\;\langle X\rangle_{j}^{T}A=2V_{j}\,$, (47)
and, since $B$ is symmetric, combining eqs. (45) and (46) results in
$V_{i}-V_{j}=\frac{1}{2}\big(\langle X\rangle_{i}^{T}V^{\prime}_{j}-\langle X\rangle_{j}^{T}V^{\prime}_{i}\big).$ (48)
Due to the results from the minimization conditions, this model has four possible EW vacua, $\mathcal{N},\mathcal{N}_{s},\mathcal{N}_{p}$ and $\mathcal{N}_{sp}$, four possible charge-breaking vacua, $\mathcal{CB},\mathcal{CB}_{s},\mathcal{CB}_{p}$ and $\mathcal{CB}_{sp}$, four possible $\mathcal{CP}$-breaking vacua, $\mathcal{CP},\mathcal{CP}_{s},\mathcal{CP}_{p}$ and $\mathcal{CP}_{sp}$, and 3 neutral vacua, $\mathcal{S},\mathcal{P}$, and $\mathcal{SP}$. We define them as follows.
$\mathcal{N}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v_{1}\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v_{2}\end{pmatrix},\quad\langle S\rangle_{0}=0,\quad\langle P\rangle_{0}=0,$ (49)
$\mathcal{N}_{s}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v^{\prime}_{1}\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v^{\prime}_{2}\end{pmatrix},\quad\langle S\rangle_{0}=v^{\prime}_{S},\quad\langle P\rangle_{0}=0,$ (50)
$\mathcal{N}_{p}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v^{\prime\prime}_{1}\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v^{\prime\prime}_{2}\end{pmatrix},\quad\langle S\rangle_{0}=0,\quad\langle P\rangle_{0}=v^{\prime\prime}_{P},$ (51)
$\mathcal{N}_{sp}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v^{\prime\prime\prime}_{1}\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ v^{\prime\prime\prime}_{2}\end{pmatrix},\quad\langle S\rangle_{0}=v^{\prime\prime\prime}_{S},\quad\langle P\rangle_{0}=v^{\prime\prime\prime}_{P},$ (52)
$\mathcal{CB}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ c_{1}\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}c_{2}\\ c_{3}\end{pmatrix},\quad\langle S\rangle_{0}=0,\quad\langle P\rangle_{0}=0,$ (53)
$\mathcal{CB}_{s}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ c^{\prime}_{1}\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}c^{\prime}_{2}\\ c^{\prime}_{3}\end{pmatrix},\quad\langle S\rangle_{0}=c^{\prime}_{S},\quad\langle P\rangle_{0}=0,$ (54)
$\mathcal{CB}_{p}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ c^{\prime\prime}_{1}\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}c^{\prime\prime}_{2}\\ c^{\prime\prime}_{3}\end{pmatrix},\quad\langle S\rangle_{0}=0,\quad\langle P\rangle_{0}=c^{\prime\prime}_{P},$ (55)
$\mathcal{CB}_{sp}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ c^{\prime\prime\prime}_{1}\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}c^{\prime\prime\prime}_{2}\\ c^{\prime\prime\prime}_{3}\end{pmatrix},\quad\langle S\rangle_{0}=c^{\prime\prime\prime}_{S},\quad\langle P\rangle_{0}=c^{\prime\prime\prime}_{P},$ (56)
$\mathcal{CP}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ \bar{v}_{1}\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ \bar{v}_{2}+i\bar{v}_{3}\end{pmatrix},\quad\langle S\rangle_{0}=0,\quad\langle P\rangle_{0}=0,$ (57)
$\mathcal{CP}_{s}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ \bar{v}^{\prime}_{1}\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ \bar{v}^{\prime}_{2}+i\bar{v}^{\prime}_{3}\end{pmatrix},\quad\langle S\rangle_{0}=\bar{v}^{\prime}_{S},\quad\langle P\rangle_{0}=0,$ (58)
$\mathcal{CP}_{p}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ \bar{v}^{\prime\prime}_{1}\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ \bar{v}^{\prime\prime}_{2}+i\bar{v}^{\prime\prime}_{3}\end{pmatrix},\quad\langle S\rangle_{0}=0,\quad\langle P\rangle_{0}=\bar{v}^{\prime\prime}_{P},$ (59)
$\mathcal{CP}_{sp}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ \bar{v}^{\prime\prime\prime}_{1}\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ \bar{v}^{\prime\prime\prime}_{2}+i\bar{v}^{\prime\prime\prime}_{3}\end{pmatrix},\quad\langle S\rangle_{0}=\bar{v}^{\prime\prime\prime}_{S},\quad\langle P\rangle_{0}=\bar{v}^{\prime\prime\prime}_{P},$ (60)
$\mathcal{S}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ 0\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ 0\end{pmatrix},\quad\langle S\rangle_{0}=S,\quad\langle P\rangle_{0}=0,$ (61)
$\mathcal{P}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ 0\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ 0\end{pmatrix},\quad\langle S\rangle_{0}=0,\quad\langle P\rangle_{0}=P,$ (62)
$\mathcal{SP}\to\langle\Phi_{1}\rangle_{0}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ 0\end{pmatrix},\quad\langle\Phi_{2}\rangle_{0}=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ 0\end{pmatrix},\quad\langle S\rangle_{0}=S,\quad\langle P\rangle_{0}=P.$ (63)
The vacuum in which we focus our analysis is the $\mathcal{N}$-type vacuum in eq. (49).
4 Sufficient BFB conditions
We now will use a parameterization inspired by Ref. Faro and Ivanov (2019),
$\Phi_{1}=\sqrt{r_{1}}\begin{pmatrix}0\\ 1\end{pmatrix},\quad\Phi_{2}=\sqrt{r_{2}}\begin{pmatrix}\sin(\alpha_{2})\\ \cos(\alpha_{2})e^{i\beta_{2}}\end{pmatrix},\quad S=\eta_{S}\sqrt{r_{3}},\quad P=\eta_{P}\sqrt{r_{4}}\,$, (64)
where $\eta_{S,P}=\pm 1$ and $r_{i}\geq 0$. With this parameterization we can show that $V_{N}$ can be written as a quadratic form,
$V_{N}=\frac{1}{2}\sum_{ij}r_{i}A_{ij}r_{j}\,$, (65)
with the matrix $A$ given by
$A=\begin{bmatrix}\lambda_{1}&\lambda_{3}+\lambda_{4}&\frac{1}{2}\lambda_{7}&\frac{1}{2}\lambda_{11}\\ \lambda_{3}+\lambda_{4}&\lambda_{2}&\frac{1}{2}\lambda_{8}&\frac{1}{2}\lambda_{12}\\ \frac{1}{2}\lambda_{7}&\frac{1}{2}\lambda_{8}&\frac{1}{4}\lambda_{6}&\frac{1}{4}\lambda_{10}\\ \frac{1}{2}\lambda_{11}&\frac{1}{2}\lambda_{12}&\frac{1}{4}\lambda_{10}&\frac{1}{4}\lambda_{9}\end{bmatrix}\,$. (66)
This part of the potential is BFB if this form is positive definite for $r_{i}\geq 0$.
The problem is that we have to find the conditions for a matrix of order four to be definite positive. For the case of matrices of order three there is a simple result. A $3\times 3$ symmetric matrix is positive definite if and only if the following conditions, known as copositivity conditions Klimenko (1985); Kannike (2012), are satisfied:
$A_{11}\geq 0,A_{22}\geq 0,A_{33}\geq 0\,$,
$\overline{A}_{12}=\sqrt{A_{11}A_{22}}+A_{12}\geq 0,\quad\overline{A}_{13}=\sqrt{A_{11}A_{33}}+A_{13}\geq 0,\quad\overline{A}_{23}=\sqrt{A_{22}A_{33}}+A_{23}\geq 0\,$,
$\sqrt{A_{11}A_{22}A_{33}}+A_{12}\sqrt{A_{33}}+A_{13}\sqrt{A_{22}}+A_{23}\sqrt{A_{11}}+\sqrt{2\overline{A}_{12}\overline{A}_{13}\overline{A}_{23}}\geq 0.$ (67)
We will come back to these expressions after we find a quadratic form that bounds the potential from below. As the odd part, $V_{O}$, cannot be written as a quadratic form, one has to find sufficient although not necessary conditions for the BFB Boto et al. (2022). For this, we bound each part of the potential. We have
$V_{CB}\geq V_{CB}^{\rm lower}=r_{1}r_{2}\,\text{min}(0,-\lambda_{4})\,$, (68)
where we have used
$0\leq z_{12}\leq 1\,.$ (69)
For $V_{HC}$ we get
$V_{HC}\geq V_{HC}^{\rm lower}=-|\lambda_{5}|r_{1}r_{2}\,.$ (70)
Finally the odd part, $V_{O}$ is the more complicated. We have
$V_{O}=\frac{1}{6}\lambda_{13}r_{3}\eta_{S}\eta_{P}\sqrt{r_{3}}\sqrt{r_{4}}+\frac{1}{6}\lambda_{14}r_{4}\eta_{S}\eta_{P}\sqrt{r_{3}}\sqrt{r_{4}}+\frac{1}{2}\left(\lambda_{15}r_{1}+\lambda_{16}r_{2}\right)\eta_{S}\eta_{P}\sqrt{r_{3}}\sqrt{r_{4}}\,.$ (71)
Let us consider the first term. As all the $r_{i}$ are positive definite, the worst situation occurs when $\lambda_{13}\eta_{S}\eta_{P}=-|\lambda_{13}|$. Then certainly we have,
$V_{O}^{\rm 1st}\geq-\frac{1}{6}\,|\lambda_{13}|r_{3}\sqrt{r_{3}}\sqrt{r_{4}}\;.$ (72)
Now we use the relation (for $r_{i}\geq 0$)
$\sqrt{r_{3}}\sqrt{r_{4}}\geq-r_{3}-r_{4}\;,$ (73)
to obtain
$V_{O}^{\rm 1st}\geq-\frac{1}{6}\,|\lambda_{13}|r_{3}(r_{3}+r_{4})\;,$ (74)
so we can safely say that
$V_{O}^{\rm 1st}\geq-\frac{1}{6}\,|\lambda_{13}|(r_{3}^{3}+r_{3}r_{4})\;,$ (75)
which is a quadratic form. Continuing with this reasoning, we obtain
$V_{O}\geq$ $V_{O}^{\rm lower}=-\frac{1}{6}\,|\lambda_{13}|(r_{3}^{2}+r_{3}r_{4})-\frac{1}{6}\,|\lambda_{14}|(r_{4}^{2}+r_{3}r_{4})$
$-\frac{1}{2}\,|\lambda_{15}|(r_{1}r_{3}+r_{1}r_{4})-\frac{1}{2}\,|\lambda_{16}|(r_{2}r_{3}+r_{2}r_{4}).$ (76)
Putting everything together, the sufficient conditions for the quadratic potential to be BFB are equivalent to requiring that the quadratic form
$\frac{1}{2}\sum_{ij}r_{i}\overline{A}_{ij}r_{j},\quad\text{where}\quad\overline{A}=\begin{bmatrix}\bar{\lambda}_{11}&\bar{\lambda}_{12}&\bar{\lambda}_{13}&\bar{\lambda}_{14}\\ \bar{\lambda}_{12}&\bar{\lambda}_{22}&\bar{\lambda}_{23}&\bar{\lambda}_{24}\\ \bar{\lambda}_{13}&\bar{\lambda}_{23}&\bar{\lambda}_{33}&\bar{\lambda}_{34}\\ \bar{\lambda}_{14}&\bar{\lambda}_{24}&\bar{\lambda}_{34}&\bar{\lambda}_{44}\end{bmatrix}$, (77)
and
$\bar{\lambda}_{11}=$ $\lambda_{1}\,$,
$\bar{\lambda}_{12}=$ $\lambda_{3}+\lambda_{4}+\text{min}(0,-\lambda_{4})-|\lambda_{5}|\,$,
$\bar{\lambda}_{13}=$ $\frac{1}{2}\lambda_{7}-\frac{1}{2}|\lambda_{15}|\,$,
$\bar{\lambda}_{14}=$ $\frac{1}{2}\lambda_{11}-\frac{1}{2}|\lambda_{15}|\,$,
$\bar{\lambda}_{22}=$ $\lambda_{2}$
$\bar{\lambda}_{23}=$ $\frac{1}{2}\lambda_{8}-\frac{1}{2}|\lambda_{16}|\,$,
$\bar{\lambda}_{24}=$ $\frac{1}{2}\lambda_{12}-\frac{1}{2}|\lambda_{16}|\,$,
$\bar{\lambda}_{33}=$ $\frac{1}{4}\lambda_{6}-\frac{1}{3}|\lambda_{13}|\,$,
$\bar{\lambda}_{34}=$ $\frac{1}{4}\lambda_{10}-\frac{1}{6}|\lambda_{13}|-\frac{1}{6}|\lambda_{14}|\,$,
$\bar{\lambda}_{44}=$ $\frac{1}{4}\lambda_{9}-\frac{1}{3}|\lambda_{14}|\,$, (78)
is positive definite. The criteria for verifying the copositivity of the $4\times 4$ matrix $\bar{A}$ are given in Ref. Ping and Yu (1993). We have implemented this long algorithm (it has 49 steps) and verified, using the CERN library Minuit James and Roos (1975), that all points that passed the above criteria were indeed BFB.
5 Global minimum
It is not only necessary to ensure that the potential is bounded from below but also to ensure that the minimum that we want to study is indeed the global minimum. This is a more demanding task because there are many other minima, and for some of them it is not possible to express analytically the value of the potential at the minimum in terms of the parameters of the potential. Just to fix notation we consider, in general,
$\left\langle\Phi_{1}\right\rangle=\begin{pmatrix}0\\ \frac{1}{\sqrt{2}}v_{1}\end{pmatrix},\ \left\langle\Phi_{2}\right\rangle=\begin{pmatrix}0\\ \frac{1}{\sqrt{2}}v_{2}\end{pmatrix},\ \left\langle S\right\rangle=v_{S},\ \left\langle P\right\rangle=v_{P}\,.$ (79)
As an example of a difficult situation is the case, $v_{1}=v_{2}=0$. The minimization conditions lead to a system of coupled non-linear equations with no simple analytical solution. However there are a few that are simple to solve and we will give here their expressions.
5.1 $v_{1}\not=0,v_{2}\not=0$ and $v_{S}=v_{P}=0$
This is the case that we want to study. It is therefore important to have an expression for this case so that we can compare with other minima. We have
$V_{N}=\frac{1}{8}\left(-\lambda_{1}v_{1}^{4}-v_{2}^{2}(2\lambda_{3}v_{1}^{2}+2\lambda_{4}v_{1}^{2}+2\lambda_{5}v_{1}^{2}+\lambda_{2}v_{2}^{2})\right)\,.$ (80)
5.2 $v_{1}=0,v_{2}=0,v_{S}=0$ and $v_{P}\not=0$
We have from the minimization equation [^2] We do not check if it is a minimum or a saddle point. This is because, if it is a saddle point, there will be a minimum below that, and therefore the point should be discarded.
$v_{P}^{2}=-\frac{2m_{P}^{2}}{\lambda_{9}}\,.$ (81)
Then, if $v_{P}^{2}>0$, the potential at he the minimum is given by,
$V_{p}=-\frac{m_{P}^{2}}{2\lambda_{9}}\,.$ (82)
5.3 $v_{1}=0,v_{2}=0,v_{P}=0$ and $v_{S}\not=0$
We have from the minimization equation
$v_{S}^{2}=-\frac{2m_{S}^{2}}{\lambda_{6}}\,.$ (83)
Then, if $v_{S}^{2}>0$, the potential at the minimum is given by,
$V_{s}=-\frac{m_{S}^{2}}{2\lambda_{6}}\,.$ (84)
5.4 $v_{1}=0,v_{S}=0$ and $v_{2}\not=0,v_{P}\not=0$
We have from the minimization equation
$v_{P}^{2}=$ $-\frac{2(2\lambda_{12}m_{22}^{2}-\lambda_{2}m_{P}^{2})}{\lambda_{12}^{2}-\lambda_{2}\lambda_{9}}\,$,
$v_{2}^{2}=$ $\frac{2(2\lambda_{9}m_{22}^{2}-\lambda_{12}m_{P}^{2})}{\lambda_{12}^{2}-\lambda_{2}\lambda_{9}}\,.$ (85)
Then, if $v_{P}^{2}>0$ and $v_{2}^{2}>0$, the potential at the minimum is given by,
$V_{v_{2}p}=\frac{4\lambda_{9}m_{22}^{4}-4\lambda_{12}m_{22}^{2}m_{P}^{2}+\lambda_{2}m_{P}^{4}}{2\lambda_{12}^{2}-2\lambda_{2}\lambda_{9}}\,.$ (86)
5.5 $v_{1}=0,v_{P}=0$ and $v_{2}\not=0,v_{S}\not=0$
We have from the minimization equation
$v_{S}^{2}=$ $-\frac{2(-2\lambda_{8}m_{22}^{2}+\lambda_{2}m_{S}^{2})}{\lambda_{2}\lambda_{6}-\lambda_{8}^{2}}\,$,
$v_{2}^{2}=$ $\frac{2(-2\lambda_{6}m_{22}^{2}+\lambda_{8}m_{S}^{2})}{\lambda_{2}\lambda_{6}-\lambda_{8}^{2}}\,.$ (87)
Then, if $v_{S}^{2}>0$ and $v_{2}^{2}>0$, the potential at the minimum is given by,
$V_{v_{2}s}=-\frac{4\lambda_{6}m_{22}^{4}-4\lambda_{8}m_{22}^{2}m_{S}^{2}+\lambda_{2}m_{S}^{4}}{2\lambda_{2}\lambda_{6}-2\lambda_{8}^{2}}\,.$ (88)
5.6 $v_{2}=0,v_{S}=0$ and $v_{1}\not=0,v_{P}\not=0$
We have from the minimization equation
$v_{P}^{2}=$ $-\frac{2(2\lambda_{11}m_{11}^{2}-\lambda_{1}m_{P}^{2})}{\lambda_{11}^{2}-\lambda_{1}\lambda_{9}}\,$,
$v_{1}^{2}=$ $\frac{2(-2\lambda_{9}m_{11}^{2}+\lambda_{11}m_{P}^{2})}{-\lambda_{11}^{2}+\lambda_{1}\lambda_{9}}\,.$ (89)
Then, if $v_{P}^{2}>0$ and $v_{1}^{2}>0$, the potential at the minimum is given by,
$V_{v_{1}p}=\frac{4\lambda_{9}m_{11}^{4}-4\lambda_{11}m_{11}^{2}m_{P}^{2}+\lambda_{1}m_{P}^{4}}{2\lambda_{11}^{2}-2\lambda_{1}\lambda_{9}}\,.$ (90)
5.7 $v_{2}=0,v_{P}=0$ and $v_{1}\not=0,v_{S}\not=0$
We have from the minimization equation
$v_{S}^{2}=$ $-\frac{2(-2\lambda_{7}m_{11}^{2}+\lambda_{1}m_{S}^{2})}{\lambda_{1}\lambda_{6}-\lambda_{7}^{2}}\,$,
$v_{1}^{2}=$ $\frac{2(-2\lambda_{6}m_{11}^{2}+\lambda_{7}m_{S}^{2})}{\lambda_{1}\lambda_{6}-\lambda_{7}^{2}}\,.$ (91)
Then, if $v_{S}^{2}>0$ and $v_{1}^{2}>0$, the potential at the minimum is given by,
$V_{v_{1}s}=-\frac{4\lambda_{6}m_{11}^{4}-4\lambda_{7}m_{11}^{2}m_{S}^{2}+\lambda_{1}m_{S}^{4}}{2\lambda_{1}\lambda_{6}-2\lambda_{7}^{2}}\,.$ (92)
5.8 $v_{2}=0,v_{P}=0,v_{S}=0$ and $v_{1}\not=0$
We have from the minimization equation
$v_{1}^{2}=$ $-\frac{4m_{11}^{2}}{\lambda_{1}}\,.$ (93)
Then, if $v_{1}^{2}>0$, the potential at the minimum is given by,
$V_{v_{1}}=-\frac{2m_{11}^{4}}{\lambda_{1}}\,.$ (94)
5.9 $v_{1}=0,v_{P}=0,v_{S}=0$ and $v_{2}\not=0$
We have from the minimization equation
$v_{2}^{2}=$ $-\frac{4m_{22}^{2}}{\lambda_{2}}\,.$ (95)
Then, if $v_{2}^{2}>0$, the potential at the minimum is given by,
$V_{v_{2}}=-\frac{2m_{22}^{4}}{\lambda_{2}}\,.$ (96)
5.10 Other minima
So, if our minimum, in eq. (80), is above one of the cases in eqs. (82),(84), (86), (88), (90), (92), (94),(96) then the point in parameter space must be discarded. However, the list above does not contain all the possible minima. We have verified with a minimization procedure using Minuit James and Roos (1975) that there are other two cases. These do not have an analytical solution (system of coupling cubic non-linear equations), but we have found a way of solving them numerically. We will discuss these in the next two subsections.
5.11 $v_{S}=v_{P}=0$, $v_{1},v_{2}\not=0$, and $\alpha_{2}=\pi$
The potential reads
$V_{\alpha_{2}=\pi}=\frac{1}{2}\left(m_{11}^{2}v_{1}^{2}+2m_{12}^{2}v_{1}v_{2}+m_{22}^{2}v_{2}^{2}\right)+\frac{1}{24}\left(3\lambda_{1}v_{1}^{4}+3\lambda_{2}v_{2}^{4}+6\lambda_{345}v_{1}^{2}v_{2}^{2}\right)\,$, (97)
where
$\lambda_{345}=\lambda_{3}+\lambda_{4}+\lambda_{5}\,.$ (98)
This leads to the minimization equations
$2\frac{\partial V}{\partial v_{1}}=$  $\,2m_{11}^{2}v_{1}+\lambda_{1}v_{1}^{3}+v_{2}(2m_{12}^{2}+\lambda_{345}v_{1}v_{2})=0\,$, (99)
$2\frac{\partial V}{\partial v_{2}}=$  $\,2m_{12}^{2}v_{1}+2m_{22}^{2}v_{2}+\lambda_{345}v_{1}^{2}v_{2}+\lambda_{2}v_{2}^{3}=0\,.$ (100)
This is a system of cubic non-linear equations with no analytical solution. However we devised a way of solving it numerically. The idea is to define
$v_{1}=v\cos\beta,\ v_{2}=v\sin\beta\,$, (101)
where $v,\beta$ should not be confused with the quantities of the same name for our minimum. Substituting and rearranging we get two equations,
$0=$  $\,2\lambda_{1}m_{12}^{2}\cos^{2}\beta-2\lambda_{345}m_{11}^{2}\cos\beta\sin\beta+2\lambda_{1}m_{22}^{2}\cos\beta\sin\beta-2\lambda_{2}m_{11}^{2}\sin^{2}\beta\tan\beta$
$+2\lambda_{345}m_{22}^{2}\sin^{2}\beta\tan\beta-2\lambda_{2}m_{12}^{2}\sin^{2}\beta\tan^{2}\beta\,$, (102)
$v^{2}=$ $\,\frac{-2m_{11}^{2}-2m_{12}^{2}\tan\beta}{\lambda_{1}\cos^{2}\beta+\lambda_{345}\sin^{2}\beta}\,.$ (103)
Now we know that with our choices $\beta\in[0,\pi/2]$. So we solve numerically the first equation for $\beta$ in that interval. If it has a solution, we substitute it in the second equation. If $v^{2}>0$ we substitute in eq. (101) and then in eq. (97). We finally compare this value with the value of the inert minimum, eq. (80). If it is lower, we discard the point.
5.12 $v_{1}=v_{2}=0$, $v_{S},v_{P}\not=0$
Finally we consider the last case, $v_{1}=v_{2}=0$, $v_{S},v_{P}\not=0$. We adopt a strategy similar to that of the last case. First the potential reduces to
$V_{s,p}=$ $\frac{1}{2}\left(m_{P}^{2}\ v_{P}^{2}-2m_{SP}\ v_{P}\,v_{S}+m_{S}^{2}\ v_{S}^{2}\right)$
$+\frac{1}{24}\left(6\lambda_{10}\ v_{P}^{2}\,v_{S}^{2}+4\lambda_{13}\ v_{P}\,v_{S}^{3}+4\lambda_{14}\ v_{P}^{3}\,v_{S}+3\lambda_{6}\ v_{S}^{4}+3\lambda_{9}\ v_{P}^{4}\right)\ .$ (104)
The stationary equations read
$\frac{\partial V}{\partial v_{S}}=$ $\frac{1}{2}\lambda_{10}\,v_{P}^{2}v_{S}+\frac{1}{2}\lambda_{13}\,v_{P}\,v_{S}^{2}+\frac{\lambda_{14}\,v_{P}^{3}}{6}+\frac{\lambda_{6}\,v_{S}^{3}}{2}-m_{SP}^{2}\,v_{P}+m_{S}^{2}\,v_{S}=0\ ,$ (105)
$\frac{\partial V}{\partial v_{P}}=$ $\frac{1}{2}\lambda_{10}\,v_{P}\,v_{S}^{2}+\frac{\lambda_{13}\,v_{S}^{3}}{6}+\frac{1}{2}\lambda_{14}\,v_{P}^{2}v_{S}+\frac{\lambda_{9}\,v_{P}^{3}}{2}+m_{P}^{2}\,v_{P}-m_{SP}^{2}\,v_{S}=0\ .$ (106)
Now we define
$v_{P}=v_{S}\,\eta\ ,$ (107)
and then eq. (105) can be solved for the following two equations,
$0=$ $\frac{1}{6}\eta^{4}\lambda_{14}m_{P}^{2}+\frac{1}{2}\eta^{4}\lambda_{9}m_{SP}^{2}+\frac{1}{2}\eta^{3}\lambda_{10}m_{P}^{2}+\frac{1}{3}\eta^{3}\lambda_{14}m_{SP}^{2}-\frac{1}{2}\eta^{3}\lambda_{9}m_{S}^{2}+\frac{1}{2}\eta^{2}\lambda_{13}m_{P}^{2}$
$-\frac{1}{2}\eta^{2}\lambda_{14}m_{S}^{2}-\frac{\eta\lambda_{10}m_{S}^{2}}{2}-\frac{\eta\lambda_{13}m_{SP}^{2}}{3}+\frac{\eta\lambda_{6}m_{P}^{2}}{2}-\frac{\lambda_{13}m_{S}^{2}}{6}-\frac{\lambda_{6}m_{SP}^{2}}{2}\ ,$ (108)
for $\eta$ and
$v_{S}^{2}=\frac{6\eta\,m_{SP}^{2}-6m_{S}^{2}}{\eta^{3}\,\lambda_{14}+3\eta^{2}\,\lambda_{10}+3\eta\,\lambda_{13}+3\lambda_{6}}\ ,$ (109)
for $v_{S}^{2}$. The procedure follows as before. We solve numerically eq. (5.12). Then substitute in eq. (109). If $v_{S}^{2}>0$, then we get the value of the potential at this stationary point substituting back in eq. (105), after using eq. (107). Then we compare with our minimum, eq. (80), and discard the point if this mininum is lower than our desired minimum. There is only one subtle point here. In principle $v_{S},v_{P}$ can have either sign, which means that also $\eta$ can be positive or negative. So we have to find solutions for $\eta$ in both cases. Then we proceed to check if $v_{S}^{2}>0$. If this is the case we still have the posibility of $v_{S}=\pm\sqrt{v_{S}^{2}}$. So we have, in principle, four possibilities, two signs for $\eta$ and two signs for $v_{S}$. We calculate the value of the potential in eq. (5.12) for all possibilities, and compare with our desired minimum, eq. (80).
6 Perturbative Unitarity
For the perturbative unitarity, we follow section 4 of Ref. Bento et al. (2017). To be self-contained we reproduce here their table 3 for our case. For convenience we define
$s_{1}=s,\quad s_{2}=p\,$, (110)
and we get the results in table 1. Notice that in comparison with Table 3 of Ref. Bento et al. (2017), we have
$Q$ $2\mathcal{Y}$ State # states
2 2 $S_{\alpha}^{++}=\{w_{1}^{+}w_{1}^{+},w_{1}^{+}w_{2}^{+},w_{2}^{+}w_{2}^{+}\}$ 33
1 2 $S_{\alpha}^{+}=\{w_{1}^{+}n_{1},w_{1}^{+}n_{2},w_{2}^{+}n_{1},w_{2}^{+}n_{2}\}$ 44
1 1 $R_{\alpha}^{+}=\{w_{1}^{+}s_{1},w_{1}^{+}s_{2},w_{2}^{+}s_{1},w_{2}^{+}s_{2}\}$ 44
1 0 $T_{\alpha}^{+}=\{w_{1}^{+}n^{*}_{1},w_{1}^{+}n^{*}_{2},w_{2}^{+}n^{*}_{1},w_{2}^{+}n^{*}_{2}\}$ 44
0 2 $S_{\alpha}^{0}=\{n_{1}n_{1},n_{1}n_{2},n_{2}n_{2}\}$ 33
0 1 $P_{\alpha}^{0}=\{n_{1}s_{1},n_{1}s_{2},n_{2}s_{1},n_{2}s_{2}\}$ 44
0 0 $T_{\alpha}^{0}=\{w_{1}^{-}w_{1}^{+},w_{1}^{-}w_{2}^{+},w_{2}^{-}w_{1}^{+},w_{2}^{-}w_{2}^{+},n_{1}n^{*}_{1},n_{1}n^{*}_{2},n_{2}n^{*}_{1},n_{2}n^{*}_{2},s_{1}s_{1},s_{1}s_{2},s_{2}s_{2}\}$ $11$
Table 1: List of two body scalar states separated by $(Q,\mathcal{Y})$.
two more possibilities, that we denote by $R^{+}_{\alpha}$ with charge 1 and hypercharge 1, and $P_{\alpha}^{0}$ with charge 0 and hypercharge 1. These are not present in NHDM. We have therefore to obtain seven matrices. From these four are equal to the 2HDM one is different and two are new.
Applying the procedure described in Ref. Bento et al. (2017) we get the following results with the notation $M_{\mathcal{Y}}^{\mathcal{Q}}$.
$M_{2}^{++}=\begin{bmatrix}\lambda_{1}&0&\lambda_{5}\\ 0&\lambda_{3}+\lambda_{4}&0\\ \lambda_{5}&0&\lambda_{2}\end{bmatrix}\,$, (111)
with eigenvalues
$\lambda_{3}+\lambda_{4},\ \frac{1}{2}(\lambda_{1}+\lambda_{2}\pm\sqrt{\lambda_{1}^{2}-2\lambda_{1}\lambda_{2}+\lambda_{2}^{2}+4\lambda_{5}^{2}})\,.$ (112)
$M_{2}^{+}=\begin{bmatrix}\lambda_{1}&0&0&\lambda_{5}\\ 0&\lambda_{3}&\lambda_{4}&0\\ 0&\lambda_{4}&\lambda_{3}&0\\ \lambda_{5}&0&0&\lambda_{2}\end{bmatrix}\,$, (113)
with eigenvalues
$\lambda_{3}-\lambda_{4},\lambda_{3}+\lambda_{4},\frac{1}{2}\left(\lambda_{1}+\lambda_{2}\pm\sqrt{\lambda_{1}^{2}-2\lambda_{1}\lambda_{2}+\lambda_{2}^{2}+4\lambda_{5}^{2}}\right)\,.$ (114)
$M_{1}^{+}=\begin{bmatrix}\lambda_{7}&\frac{1}{2}\lambda_{15}&0&0\\ \frac{1}{2}\lambda_{15}&\lambda_{11}&0&0\\ 0&0&\lambda_{8}&\frac{1}{2}\lambda_{16}\\ 0&0&\frac{1}{2}\lambda_{16}&\lambda_{12}\end{bmatrix}\,$, (115)
with eigenvalues
$\frac{1}{2}(\lambda_{11}+\lambda_{7}\pm\sqrt{\lambda_{11}^{2}+\lambda_{15}^{2}-2\lambda_{11}\lambda_{7}+\lambda_{7}^{2}}),\frac{1}{2}(\lambda_{12}+\lambda_{8}\pm\sqrt{\lambda_{12}^{2}+\lambda_{16}^{2}-2\lambda_{12}\lambda_{8}+\lambda_{8}^{2}})\,.$ (116)
$M_{0}^{+}=\begin{bmatrix}\lambda_{1}&0&0&\lambda_{4}\\ 0&\lambda_{3}&\lambda_{5}&0\\ 0&\lambda_{5}&\lambda_{3}&0\\ \lambda_{4}&0&0&\lambda_{2}\\ \end{bmatrix}\,$, (117)
with eigenvalues,
$\lambda_{3}-\lambda_{5},\lambda_{3}+\lambda_{5},\frac{1}{2}(\lambda_{1}+\lambda_{2}\pm\sqrt{\lambda_{1}^{2}-2\lambda_{1}\lambda_{2}+\lambda_{2}^{2}+4\lambda_{4}^{2}})\,.$ (118)
$M_{2}^{0}=M_{2}^{++}\,$, (119)
with the same eigenvalues as in eq. (112).
$M_{1}^{0}=M_{1}^{+}\,$, (120)
with the same eigenvalues as in eq. (116). Finally
$M_{0}^{0}=\left[\begin{array}[]{ccccccccccc}2\lambda_{1}&0&0&\lambda_{3}+\lambda_{4}&\lambda_{1}&0&0&\lambda_{3}&\frac{\lambda_{7}}{\sqrt{2}}&\frac{\lambda_{15}}{2}&\frac{\lambda_{11}}{\sqrt{2}}\\ 0&\lambda_{3}+\lambda_{4}&2\lambda_{5}&0&0&\lambda_{4}&\lambda_{5}&0&0&0&0\\ 0&2\lambda_{5}&\lambda_{3}+\lambda_{4}&0&0&\lambda_{5}&\lambda_{4}&0&0&0&0\\ \lambda_{3}+\lambda_{4}&0&0&2\lambda_{2}&\lambda_{3}&0&0&\lambda_{2}&\frac{\lambda_{8}}{\sqrt{2}}&\frac{\lambda_{16}}{2}&\frac{\lambda_{12}}{\sqrt{2}}\\ \lambda_{1}&0&0&\lambda_{3}&2\lambda_{1}&0&0&\lambda_{3}+\lambda_{4}&\frac{\lambda_{7}}{\sqrt{2}}&\frac{\lambda_{15}}{2}&\frac{\lambda_{11}}{\sqrt{2}}\\ 0&\lambda_{4}&\lambda_{5}&0&0&\lambda_{3}+\lambda_{4}&2\lambda_{5}&0&0&0&0\\ 0&\lambda_{5}&\lambda_{4}&0&0&2\lambda_{5}&\lambda_{3}+\lambda_{4}&0&0&0&0\\ \lambda_{3}&0&0&\lambda_{2}&\lambda_{3}+\lambda_{4}&0&0&2\lambda_{2}&\frac{\lambda_{8}}{\sqrt{2}}&\frac{\lambda_{16}}{2}&\frac{\lambda_{12}}{\sqrt{2}}\\ \frac{\lambda_{7}}{\sqrt{2}}&0&0&\frac{\lambda_{8}}{\sqrt{2}}&\frac{\lambda_{7}}{\sqrt{2}}&0&0&\frac{\lambda_{8}}{\sqrt{2}}&\frac{3\lambda_{6}}{2}&\frac{\lambda_{13}}{\sqrt{2}}&\frac{\lambda_{10}}{2}\\ \frac{\lambda_{15}}{2}&0&0&\frac{\lambda_{16}}{2}&\frac{\lambda_{15}}{2}&0&0&\frac{\lambda_{16}}{2}&\frac{\lambda_{13}}{\sqrt{2}}&\lambda_{10}&\frac{\lambda_{14}}{\sqrt{2}}\\ \frac{\lambda_{11}}{\sqrt{2}}&0&0&\frac{\lambda_{12}}{\sqrt{2}}&\frac{\lambda_{11}}{\sqrt{2}}&0&0&\frac{\lambda_{12}}{\sqrt{2}}&\frac{\lambda_{10}}{2}&\frac{\lambda_{14}}{\sqrt{2}}&\frac{3\lambda_{9}}{2}\end{array}\right].$ (121)
The eigenvalues separate in six that can be easily evaluated, but the other five are solutions of a fifth order polynomial. In this case it is easier to use the methods of the minors as explained in Ref. Bento et al. (2022) for the whole matrix.
7 The precision observables $S,T$ and $U$
To obtain these observables one follows the setup of Ref. Grimus et al. (2008). For this one has to obtain the matrices $U$ and $V$ in their notation.
Following Ref. Grimus et al. (2008), and using the definitions given in section 2 we define
$\begin{bmatrix}\varphi_{1}^{0}\\ \varphi_{2}^{0}\end{bmatrix}=\begin{bmatrix}y_{1}+iz_{1}\\ y_{2}+iz_{2}\end{bmatrix},\qquad\begin{bmatrix}\chi_{1}^{0}\\ \chi_{2}^{0}\end{bmatrix}=\begin{bmatrix}s\\ p\end{bmatrix}$. (122)
We have now everything to define the matrices $U_{2\times 2}$, $V_{2\times 6}$ and $R_{2\times 6}.$ [^3] As can be seen from the expressions for $S,T,U$ in Ref. Grimus et al. (2008), the matrix $R$ is not necessary. We get,
$U=\mathcal{O}_{\beta}^{T}=\begin{bmatrix}c_{\beta}&-s_{\beta}\\ s_{\beta}&c_{\beta}\end{bmatrix}$, (123)
$V=\begin{bmatrix}i\,c_{\beta}&c_{\alpha_{1}}&s_{\alpha_{1}}&-i\,s_{\beta}&0&0\\ i\,s_{\beta}&-s_{\alpha_{1}}&c_{\alpha_{1}}&i\,c_{\beta}&0&0\end{bmatrix}$, (124)
$R=\begin{bmatrix}0&0&0&0&c_{\alpha_{2}}&s_{\alpha_{2}}\\ 0&0&0&0&-s_{\alpha_{2}}&c_{\alpha_{2}}\end{bmatrix}$, (125)
where we have organized the physical fields in the order
$(G_{0},h_{1},h_{2},A,h_{3},h_{4})^{T}\,.$ (126)
The matrices required for the calculation of the precision observables are,
$U^{\dagger}U=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$, (127)
$U^{\dagger}V=\left[\begin{array}[]{cccccc}i&c_{\alpha_{1}}c_{\beta}-s_{\alpha_{1}}s_{\beta}&c_{\alpha_{1}}s_{\beta}+c_{\beta}s_{\alpha_{1}}&0&0&0\\ 0&-c_{\alpha_{1}}s_{\beta}-c_{\beta}s_{\alpha_{1}}&c_{\alpha_{1}}c_{\beta}-s_{\alpha_{1}}s_{\beta}&i&0&0\\ \end{array}\right]\,$, (128)
$V^{\dagger}V=\left[\begin{array}[]{cccccc}1&is_{\alpha_{1}}s_{\beta}-ic_{\alpha_{1}}c_{\beta}&-ic_{\alpha_{1}}s_{\beta}-ic_{\beta}s_{\alpha_{1}}&0&0&0\\ ic_{\alpha_{1}}c_{\beta}-is_{\alpha_{1}}s_{\beta}&1&0&-ic_{\alpha_{1}}s_{\beta}-ic_{\beta}s_{\alpha_{1}}&0&0\\ ic_{\alpha_{1}}s_{\beta}+ic_{\beta}s_{\alpha_{1}}&0&1&ic_{\alpha_{1}}c_{\beta}-is_{\alpha_{1}}s_{\beta}&0&0\\ 0&ic_{\alpha_{1}}s_{\beta}+ic_{\beta}s_{\alpha_{1}}&is_{\alpha_{1}}s_{\beta}-ic_{\alpha_{1}}c_{\beta}&1&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ \end{array}\right]\,$, (129)
$\text{Im}(V^{\dagger}V)=\left[\begin{array}[]{cccccc}0&s_{\alpha_{1}}s_{\beta}-c_{\alpha_{1}}c_{\beta}&-c_{\alpha_{1}}s_{\beta}-c_{\beta}s_{\alpha_{1}}&0&0&0\\ c_{\alpha_{1}}c_{\beta}-s_{\alpha_{1}}s_{\beta}&0&0&-c_{\alpha_{1}}s_{\beta}-c_{\beta}s_{\alpha_{1}}&0&0\\ c_{\alpha_{1}}s_{\beta}+c_{\beta}s_{\alpha_{1}}&0&0&c_{\alpha_{1}}c_{\beta}-s_{\alpha_{1}}s_{\beta}&0&0\\ 0&c_{\alpha_{1}}s_{\beta}+c_{\beta}s_{\alpha_{1}}&s_{\alpha_{1}}s_{\beta}-c_{\alpha_{1}}c_{\beta}&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ \end{array}\right].$ (130)
Using these matrices, we have programmed the expressions in Ref. Grimus et al. (2008). The only important point to remember is the order of the fields in eq. (126).
8 Experimental Constraints
We have discussed the conditions for boundedness from below, a global minimum and perturbative unitarity. We complete the theoretical constraints by requiring perturbativity of the Type-II Yukawa couplings, by setting them to be $|y_{i}|<\sqrt{4\pi}$, where $i=t,b,\tau$. The bounds from experimental collider searches follow: the oblique parameters STU described in Section 7 are to be compared with the global electroweak fit in Baak et al. (2014); the coupling-strength modifiers within $3\sigma$ of the LHC data Aad and others (2020); the LHC signal strengths of the $125\textrm{GeV}$ Higgs, for the combinations of production cross sections and branching ratios, to have a $2\sigma$ agreement with the most recent ATLAS results Aad and others (2022). For LHC searches for new particles, we use the software package HiggsTools-1.1.3 Bahl et al. (2023), which includes the latest data from the ATLAS and CMS experiments at CERN, and the latest LHC bound on the invisible branching ratio of the Higgs, BR($h\rightarrow\text{inv}$) $\leq$ 0.107 at $95\%$ C.L. Aad and others (2024).
For the DM observables, we implemented our model in micrOMEGAs-6.2.3 Alguero et al. (2024) to numerically calculate the relic density, scattering amplitudes, annihilation cross section and decay rates. We note that the $\mathbbm{Z}_{2}$ symmetry used in this work does not separate the DM particle into two non-communicating sectors. One could think that a second DM particle could be invoked by imposing a kinematic limit on the mass of the second DM candidate. However, in this case, such putative DM candidate could decay into the (true) lightest DM candidate and a SM pair, in particular through the terms proportional to $\lambda_{15}$ and $\lambda_{16}$ in Eq. (3). Such decays are included using the micrOMEGAs routine darkOmegaN but not under the darkOmega2 routine Alguero et al. (2024), which comes activated by default in the usual micrOMEGAs distribution. We compare with the Planck experiment Aghanim and others (2020) at $3\sigma$,
$\Omega h^{2}=0.1200\pm 0.0012\,.$ (131)
While optimizing our initial simulations and comparing methods, we consider a more permissive range $0.09<\Omega h^{2}<0.15$, as identified appropriately in the respective captions.
For the DM-nucleon scattering, we follow the method described in Ref. Bélanger et al. (2014) of computing the normalized cross section of DM on a point-like Xenon nucleus
$\sigma_{\text{SI}}^{\text{Xe}}=\frac{4\mu_{\text{DM}}^{2}}{\pi}\frac{\left(Zf_{p}+(A-Z)f_{n}\right)^{2}}{A^{2}},$ (132)
with $\mu_{\text{DM}}$ the reduced mass of the DM candidate and $f_{p}$, $f_{n}$ the amplitudes for protons and neutrons. We compare with the most recent LZ release in 2024 Aalbers and others (2024a).
The annihilation cross section calculated is to be compared with reconstructions based on indirect searches. We follow the method described in Belanger et al. (2022) and compare the dominating channel with the respective experimental exclusion line. We find that for our model and mass ranges, the annihilation occurs either into $VV$, summing only the $WW$ and $ZZ$ final states, dubbed $\langle\sigma v\rangle_{VV}$, or into $b\bar{b}$  [^4] There are currently no published equivalent exclusion lines for decays going mainly into $c\bar{c}$ nor $h_{1}h_{1}$.. For each point in parameter space, the procedure is to check the annihilation channel and apply the respective exclusion bounds in Fig. 1. We have confirmed that for the mass region studied, this constraint does not exclude a significant amount of the parameter space when applied after demanding the correct relic density.
Figure 1: Combined relevant limits from indirect searches on the total $\langle\sigma v\rangle$ as a function of the mass of the DM candidate $m_{DM}$. The lines coming from Fermi-LAT Ackermann and others (2015) and H.E.S.S. Abdalla and others (2022) assume a Navarro-Frenk-White (NFW) DM density profile and the AMS-02 Aguilar and others (2016) lines correspond to the conservative approach derived in Ref. Reinert and Winkler (2018).
9 Sampling Methods
[…]
$m_{H^{+}}>580\,\text{GeV}\,.$ (133)
We continued with a longer duration scan of the parameter space. Our fixed inputs are $v=246\,\text{GeV}$ and $m_{h1}=125\,\text{GeV}$. We then took random values in the ranges:
[…]
$m_{h2},\,m_{h3},\,m_{h4}\,\in\left[125,1000\right]\,\text{GeV};$ (134b)
$m_{A},\,\in\left[100,1000\right]\,\text{GeV};\qquad m_{H^{\pm}}\in\left[580,1000\right]\,\text{GeV};$ (134c)
$m^{2}_{12},m^{2}_{S},m^{2}_{P},m^{2}_{SP}\in\left[\pm 10^{-1},\pm 10^{7}\right]\,\text{GeV}^{2}\,;$ (134d)
$\lambda_{6},\lambda_{8},\lambda_{9},\lambda_{10},\lambda_{12},\lambda_{13},\lambda_{14},\lambda_{16}\in\left[\pm 10^{-3},\pm 10^{1}\right]\,.$ (134e)
For the sampling near the alignment limit, we consider the same range, with the change that $\alpha_{1}$ is obtained as a random number within $\pm 10\%$ of $\beta$. We let each method run for about $\sim 7000$ CPU hours in order to obtain $\sim 120000$ points for the random scan and $\sim 370000$ points with the near alignment consideration.
We present the results of the random scan in Fig. 2, for the $\alpha_{1}-\beta$ plane, showing clearly that considering $\alpha_{1}$ within $\pm 10\%$ of $\beta$ correctly explores the allowed parameter space, and the scan near the alignment limit in Fig. 3 for the mass - relic density plane. The points in red satisfy BFB, unitarity, global minimum, flavour bounds, coupling modifiers and signal strengths. The points in green combine points originally in red that are found to also satisfy HiggsTools-1.1.3. We add in blue the points originally in green found to have a relic density of $\Omega h^{2}\in[0.09,0.15]$.
Figure 2: Points obtained with a random sampling using the ranges of parameters in eq. (134), shown in the $\alpha_{1}-\beta$ plane. The points in red consider the expressions of Sections 3-8, in order to satisfy BFB, unitarity, global minimum, flavour bounds, coupling modifiers and signal strengths. The points in green combine points originally in red that are found to also satisfy HiggsTools-1.1.3. The blue point is a green point that, in addition, meets the condition $\Omega h^{2}\in[0.09,0.15]$.
At this stage, we do not impose the bounds from Dark Matter direct and indirect searches. Such bounds rule out most of the few points we found with acceptable relic density, with Fig.4 showing the comparison of the points sampled near the alignment limit with the most recent exclusion line from the LZ collaboration Aalbers and others (2024a). As most points sampled do not have a significant relic density, we continue with a new Machine Learning method instead of continuing with longer sets of inefficient traditional sampling.
Figure 3: Results in the mass - relic density plane for the scan near the alignment limit, considering eq. (134), except for $\alpha$ obtained as a random number within $\pm 10\%$ of $\beta$. The color code coincides with Fig. 2.
Figure 4: Direct detection results for the scan near the alignment limit $\alpha_{1}=-[0.9,1.1]\beta$, with the relevant quantity for nucleon scattering $\sigma_{SI}^{Xe}$ obtained from eq. (132). The color code is the same as Fig. 2, with the addition of the LZ exclusion line from 2024 Aalbers and others (2024a) drawn in blue.
We followed the procedure for efficiently sampling the parameter space shown in de Souza et al. (2023) and further developed in Romão and Crispim Romão (2024); de Souza et al. (2025b, a)  [^5] Alternative approaches to successfully sample the parameter space of BSM models have been developed in cases where training datasets are available Caron et al. (2019); Hollingsworth et al. (2021); Goodsell and Joury (2023); Diaz et al. (2024); Batra et al. (2025); Hammad et al. (2025). For more broad applications of Machine Learning in the field of particle physics, consider the comprehensive reviews Feickert and Nachman (2021); Plehn et al. (2022).. The first step is to define constraint functions, $C(\mathcal{O})$, as
$C(\mathcal{O})=\max(0,-\mathcal{O}+\mathcal{O}_{LB},\mathcal{O}-\mathcal{O}_{UB}),$ (135)
with $\mathcal{O}$ the value of a quantity constrained to be inside the interval $\mathcal{O}_{LB},\mathcal{O}_{UB}]$. $C(\mathcal{O})$ then quantifies how far the value of the observable is from the defined bounds, or zero if within the specified interval. The quantities $\mathcal{O}$ are obtained by a black box computational routine that takes a set of parameters $\theta$ as inputs to calculate all relevant physical quantities $\mathcal{O}(\theta)$. We follow the single-objective optimisation algorithm, obtaining the loss function as the sum of all the constraint functions of the model
$L(\theta)=\sum_{i=1}^{N_{c}}C(\mathcal{O}_{i}(\theta)),$ (136)
where the sum runs over all the $N_{c}$ constraints, with $L=0$ only when all constraints are satisfied. One of the main strengths of the method is that the quantity $\mathcal{O}_{i}$ does not need to be an experimental observable, also allowing theoretical constraints and cuts in the same loss function.
The optimization algorithm is the Covariant Matrix Adaptation Evolutionary Strategy (CMA-ES) Hansen and Ostermeier (2001); Hansen (2023), characterized by an iterative sampling according to a multivariate normal distribution, initialized with its mean at a random point in parameter space and its covariance matrix set to the identity matrix, $\mathbb{1}$, scaled by a constant. A generation of candidate solutions is sampled from this distribution and the candidates are ranked from best to worst based on the loss function in eq. (136). The best candidates are then used to compute a new mean and approximate the covariance matrix, for the next iterative step.
CMAES has one critical downside: it has a limited exploration capacity due to the highly localized nature of the algorithm. To mitigate this, similarly to Ref. Romão and Crispim Romão (2024), we implemented a novelty reward into the loss function. This is done by adding the density of found points as a penalty to the loss function, thus allowing it to be minimal when the point is good and away from other good points. To this end, we used the Histogram Based Outlier System (HBOS), which computes a density penalty between 0 and 1, such that a new point has 0 and a point similar to previous ones approach the maximum 1, basing itself on the abundance values for chosen parameters.
To ensure proper minimization with these penalties, the loss function is shifted accordingly,
$\tilde{L}(\theta)=\left\{\begin{array}[]{ll}1+L(\theta)&\text{if }L(\theta)>0\\ \text{ }\\ 0&\text{if }L(\theta)=0\end{array}\right.$. (137)
The penalties are then added to obtain the final version of the loss function,
$L_{T}(\theta)=\tilde{L}(\theta)+\frac{1}{2}\left(\frac{1}{N^{\mathcal{P}}_{p}}\sum_{i=1}^{N^{\mathcal{P}}_{p}}p^{\mathcal{P}}_{i}(\theta^{i})+\frac{1}{N^{\mathcal{O}}_{p}}\sum_{i=1}^{N^{\mathcal{O}}_{p}}p^{\mathcal{O}}_{i}[\mathcal{O}^{i}(\theta)]\right),$ (138)
where $p^{\mathcal{P}}_{i}(\theta^{i})$ is the density penalty in parameter space $\mathcal{P}$, normalized by the amount of parameter penalties considered, $N^{\mathcal{P}}_{p}$, and $p^{\mathcal{O}}_{i}[\mathcal{O}^{i}(\theta)]$ is the density penalty of the observable space $\mathcal{O}$, also properly normalized by $N^{\mathcal{O}}_{p}$. We highlight the fact that penalties do not need to apply to all parameters $\theta$ and/or observables $\mathcal{O}(\theta)$. A subset of interest can be chosen to perform focused runs with density penalty on specific parameters and/or observables.
With the default setup, each optimization run is independent, as CMA-ES is initialized with new values for the mean and parameters in the covariant matrix and trained solely on points from that run. We may however choose valid points from previous runs as seeds to start new runs with CMA-ES initialized already in that region.
We implemented the optimization with CMA-ES and performed a first simulation, including a $C(\mathcal{O})$ for the relic density, without novelty reward and not considering bounds from DM direct/indirect searches. On $\sim 1500$ CPU hours, we obtained $\sim 660000$ points satisfying $\Omega_{T}h^{2}\in[0.1164,0.1236]$, shown in Fig. 5, for the same plane as in Figs. 3 and 4. The points in red fail one of the constraints in Sections 48, identified by least one $C(\mathcal{O})$ in the loss function not zero. The points in green are found to satisfy all constraints apart from DM bounds, satisfying BFB, unitarity, global minimum, flavour bounds, coupling modifiers and HiggsTools-1.1.3. In blue are the points originally in green found to also meet the Dark Matter constraints. The iterative process quickly moves towards the regions with considerable relic density and we are able to obtain points also satisfying direct detection bounds. However, the lack of novelty reward results in a large concentration of points in the same mass region for the DM candidate, around $65$-$300\textrm{GeV}$. In the next section we follow the method with novelty reward, that considers the relevant DM bounds with appropriate $C(\mathcal{O})$ for each experimental exclusion.
Figure 5: Results from a run using the optimization algorithm CMA-ES, without the novelty reward or seeded runs methods described in the text. In the Left panel we show the rapid convergence towards the imposed interval on the total relic density, $\Omega_{T}h^{2}\in[0.1164,0.1236]$, and in the Right panel the obtained value for the direct detection quantity $\sigma_{\text{SI}}^{\text{Xe}}$ given by eq. (132), with the most recent constraint from LZ Aalbers and others (2024a), shown as the blue line. The points in red fail one of the constraints not related to DM and the points in green meet all constraints except DM. The points with a blue color are points in green that also satisfy the correct relic density and direct detection scattering bounds.
10 Final Machine Learning Results
Following the initial set of simulations showing the ability of the evolutionary algorithm in obtaining parameter points with the correct relic density, but lacking diversity in the Dark Matter mass plane, we set our goal on fully exploring the allowed parameter space adding the novelty reward technique.
When we include indirect constraints, with exclusion lines shown in Fig. 1, we observe an increase in the time per generation of a factor of 4 and an increased convergence time. The strategy followed was then to only use micrOMEGAs-6.2.3 Alguero et al. (2024) to numerically calculate the relic density and direct detection variables. For a set of simulations, we took parameter points that met the relic density from Planck and nucleon scattering constraints to check the annihilation bounds. As a next step, we passed again the valid (relic and direct detection) points through micrOMEGAs-6.2.3, probing for consistency with indirect detection bounds. We found the large majority of points to already satisfy the current indirect detection constraints. The final results, also satisfying all other constraints described in Sections 48, have the Dark Matter relic density within $3\sigma$ of the limits obtained by the Planck experiment Aghanim and others (2020), satisfy the most recent direct detection bounds from LZ Aalbers and others (2024a), as well as the indirect detection constraints from Fig. 1. All other constraints described in Sections 48 are also considered.
We start with some run, which typically will yield one of the line structures visible in Fig. 5. Next, we take a set of valid points, and we start seeded runs based on those, specifically targeting a novel mass region. For example, we looked specially into regions where the DM candidate lies below the mass of $m_{h1}=125\,\text{GeV}$ and also regions above $500~\text{GeV}$. We believe that the gaps in the mass plane can be filled completely in the mass region shown. The allowed mass region for the DM candidate spans the entire mass range from half the Higgs mass to the high GeV scale.
No valid points were found in the mass region below $m_{h_{1}}/2$, as the latest LHC bound on the invisible branching ratio of the Higgs, BR($h\rightarrow\text{inv}$) $\leq$ 0.107 at $95\%$ C.L. Aad and others (2024), comes into effect.
In Fig. 6, we show the constraints from direct detection [^6] The data files and general layout for the plots was obtained from the public repository by Ciaran O’Hare avaliable in https://github.com/cajohare/DirectDetectionPlots. for the mass region studied. The solid lines shown correspond to the most recent exclusion bounds from experiments, with XENONnT Aprile and others (2025), PandaX-4T Bo and others (2025) and LZ Aalbers and others (2024a). The dashed lines show the projections for DarkSide-20k Aalseth and others (2018) and XLZD Aalbers and others (2024b). The neutrino floor is shown in grey, as defined in Ref. O’Hare (2021).
Figure 6: The red region correponds to valid points generated with our model. The solid lines shown correspond to the most recent exclusion bounds from experiments, with XENONnT Aprile and others (2025), PandaX-4T Bo and others (2025) and LZ Aalbers and others (2024a). The dashed lines show the projections for DarkSide-20k Aalseth and others (2018) and XLZD Aalbers and others (2024b). The neutrino floor is shown in grey, as defined in Ref. O’Hare (2021).
As one can see from Fig. 6, there are many valid points obeying the current LZ bounds (in addition to matching the correct relic density). Some of those points could be excluded (or confirmed) by future experiments, such as DarkSide-20k or XLZD. However, the fact that many red (valid model) points lie within the grey (neutrino fog) region, implies that this model will be hard to exclude with future direct detection experiments  [^7] There is also the possibility of improving theoretical and experimental tools in order to break the degeneracy between neutrino backgrounds and potential WIMP signals Davis (2015); Gelmini et al. (2018); Arcadi et al. (2025); Vahsen and others (2020)..
11 Conclusion
We have studied in detail a model with a type II 2HDM supplemented by two inert scalars which provide a viable dark matter particle. We developed here for the first time a full study of: the possible vaccua; the (sufficient) bounded from below requirements; the conditions under which our chosen vacuum - $\mathcal{N}$ in eq. (49) - is the global minimum; and the perturbative unitarity constraints.
Points in parameter space where searched for, which obeyed these conditions and, in addition, satisfy current experimental bounds arising from the oblique radiative parameters, current LHC bounds on the $125\textrm{GeV}$ Higgs as well as searches for additional scalar particles, and also direct, indirect, and relic abundance constraints. These constraints have been implemented using HiggsTools-1.1.3 Bahl et al. (2023) and micrOMEGAs-6.2.3 Alguero et al. (2024).
Three different scanning strategies have been utilized: i)  a random scan without any prior assumptions on the parameter space; ii)  a scan close to the alignment limit of the 2HDM, defined as $\alpha_{1}=-\beta$ for the choices in eqs. (13) and (15); iii)  employing the Artificial Intelligence black box optimization approach. We have found that the random scan and the scan close to the alignment limit of the 2HDM produce similar results, although (of course) the latter is more efficient than the first, by a factor close to three. Neither of these produce points with the correct relic density in a reasonable time frame. Thus, we turned to a machine learning method. Using unseeded runs, we quickly find solutions with reasonable relic density, but yielding limited exploration of the various mass regions. In contrast, using a novelty reward strategy and seeded runs, we are able to find valid points in a larger whole mass region: $125\textrm{GeV}/2<m_{DM}\lesssim 1000\textrm{GeV}$.
We have found that current/projected direct detection experiments do/will not exclude this model, since there are many valid points with the neutrino fog region. We have also studied current indirect detection constraints, and found that they do not significantly affect those points already allowed by direct detection bounds and relic density.
All models addressing the DM problem with scalar DM candidates must contend with a plethora of experimental results, from collider and astrophysical searches. When models have many parameters, such searches are very demanding computationally. One can easily seem to rule out a valid model or be tempted to take as a generic model feature, properties which were merely the result of searches performed in a region where valid points are easy to produce. This work illustrates the importance of using Machine Learning coupled with novelty detection techniques in order to efficiently explore the validity and generic features of DM models where extensive collider data must also be contended with. Harnessing the power of Machine Learning and novelty detection, opens the door to deeper, more reliable insights into dark matter models—paving the way for discovery in even the most complex parameter spaces.
12 Acknowledgments
This work is supported in part by the Portuguese Fundação para a Ciência e Tecnologia (FCT) through the PRR (Recovery and Resilience Plan), within the scope of the investment "RE-C06-i06 - Science Plus Capacity Building", measure "RE-C06-i06.m02 - Reinforcement of financing for International Partnerships in Science, Technology and Innovation of the PRR", under the project with reference 2024.01362.CERN. The work of the authors is also supported by FCT under Contracts UIDB/00777/2020, and UIDP/00777/2020. The FCT projects are partially funded through POCTI (FEDER), COMPETE, QREN, and the EU. The work of R. Boto is also supported by FCT with the PhD grant PRT/BD/152268/2021.
Appendix A Explicit expressions for depth of minimum
For the desired $\mathcal{N}$-type vacuum to be stable, it must be the global minimum of the potential. To verify that, we used equation (48) to compare this vacuum with the other minima expressed in section 3.2 and arrived at the following conclusions.
A.1 $\mathcal{N}$-type vacuum
The expressions for comparing the depth of the $\mathcal{N}$-type potential against $\mathcal{CB}$ and $\mathcal{CP}$-type vacua are:
$\mathcal{V}_{\mathcal{CB}}-\mathcal{V}_{\mathcal{N}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4v^{2}}\right)_{\mathcal{N}}\big[(v_{2}c_{1}-v_{1}c_{3})^{2}+v_{1}^{2}c_{2}^{2}\big],$ (139)
$\mathcal{V}_{\mathcal{CB}_{s}}-\mathcal{V}_{\mathcal{N}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4v^{2}}\right)_{\mathcal{N}}\big[(v_{2}c^{\prime}_{1}-v_{1}c^{\prime}_{3})^{2}+v_{1}^{2}{c^{\prime}_{2}}^{2}\big]+\frac{1}{4}{c^{\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}},$ (140)
$\mathcal{V}_{\mathcal{CB}_{p}}-\mathcal{V}_{\mathcal{N}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4v^{2}}\right)_{\mathcal{N}}\big[(v_{2}c^{\prime\prime}_{1}-v_{1}c^{\prime\prime}_{3})^{2}+v_{1}^{2}{c^{\prime\prime}_{2}}^{2}\big]+\frac{1}{4}{c^{\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}},$ (141)
$\mathcal{V}_{\mathcal{CB}_{sp}}-\mathcal{V}_{\mathcal{N}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4v^{2}}\right)_{\mathcal{N}}\big[(v_{2}c^{\prime\prime\prime}_{1}-v_{1}c^{\prime\prime\prime}_{3})^{2}+v_{1}^{2}{c^{\prime\prime\prime}_{2}}^{2}\big]+\frac{1}{4}{c^{\prime\prime\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}}+\frac{1}{4}{c^{\prime\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}}$
$+\frac{1}{2}c^{\prime\prime\prime}_{S}c^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{N}},$ (142)
$\mathcal{V}_{\mathcal{CP}}-\mathcal{V}_{\mathcal{N}}=$ $\left(\frac{m_{A}^{2}}{4v^{2}}\right)_{\mathcal{N}}\big[(v_{2}\bar{v}_{1}-v_{1}\bar{v}_{2})^{2}+v_{1}^{2}\bar{v}_{3}^{2}\big],$ (143)
$\mathcal{V}_{\mathcal{CP}_{s}}-\mathcal{V}_{\mathcal{N}}=$ $\left(\frac{m_{A}^{2}}{4v^{2}}\right)_{\mathcal{N}}\big[(v_{2}\bar{v}^{\prime}_{1}-v_{1}\bar{v}^{\prime}_{2})^{2}+v_{1}^{2}{\left.\bar{v}^{\prime}_{3}\right.}^{2}\big]+\frac{1}{4}{\left.\bar{v}^{\prime}_{S}\right.}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}},$ (144)
$\mathcal{V}_{\mathcal{CP}_{p}}-\mathcal{V}_{\mathcal{N}}=$ $\left(\frac{m_{A}^{2}}{4v^{2}}\right)_{\mathcal{N}}\big[\left(v_{2}\bar{v}^{\prime\prime}_{1}-v_{1}\bar{v}^{\prime\prime}_{2}\right)^{2}+v_{1}^{2}{\left.\bar{v}^{\prime\prime}_{3}\right.}^{2}\big]+\frac{1}{4}{\left.\bar{v}^{\prime\prime}_{P}\right.}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}},$ (145)
$\mathcal{V}_{\mathcal{CP}_{sp}}-\mathcal{V}_{\mathcal{N}}=$ $\left(\frac{m_{A}^{2}}{4v^{2}}\right)_{\mathcal{N}}\big[(v_{2}\bar{v}^{\prime\prime\prime}_{1}-v_{1}\bar{v}^{\prime\prime\prime}_{2})^{2}+v_{1}^{2}{\left.\bar{v}^{\prime\prime\prime}_{3}\right.}^{2}\big]+\frac{1}{4}{\left.\bar{v}^{\prime\prime\prime}_{S}\right.}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}}+\frac{1}{4}{\left.\bar{v}^{\prime\prime\prime}_{P}\right.}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}}$
$+\frac{1}{2}\bar{v}^{\prime\prime\prime}_{S}\bar{v}^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{N}}.$ (146)
From these equations, one can infer that if the potential has a minimum of type $\mathcal{N}$, any stationary point of type $\mathcal{CB},\mathcal{CB}_{s},\mathcal{CB}_{p},\mathcal{CP},\mathcal{CP}_{s}$ or $\mathcal{CP}_{p}$, if it exists, lies above $\mathcal{N}$. This is not necessarily the case for stationary points of type $\mathcal{CB}_{sp}$ or $\mathcal{CP}_{sp}$.
A.2 $\mathcal{N}_{s}$-type vacuum
The expressions for comparing the depth of the $\mathcal{N}_{s}$-type potential against $\mathcal{CB}$ and $\mathcal{CP}$-type vacua are:
$\mathcal{V}_{\mathcal{CB}}-\mathcal{V}_{\mathcal{N}_{s}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime}}^{2}}\right)_{\mathcal{N}_{s}}\big[(v^{\prime}_{2}c_{1}-v^{\prime}_{1}c_{3})^{2}+{v^{\prime}_{1}}^{2}c_{2}^{2}\big]-\frac{1}{4}{v^{\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{CB}},$ (147)
$\mathcal{V}_{\mathcal{CB}_{s}}-\mathcal{V}_{\mathcal{N}_{s}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime}}^{2}}\right)_{\mathcal{N}_{s}}\big[(v^{\prime}_{2}c^{\prime}_{1}-v^{\prime}_{1}c^{\prime}_{3})^{2}+{v^{\prime}_{1}}^{2}{c^{\prime}_{2}}^{2}\big],$ (148)
$\mathcal{V}_{\mathcal{CB}_{p}}-\mathcal{V}_{\mathcal{N}_{s}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime}}^{2}}\right)_{\mathcal{N}_{s}}\big[(v^{\prime}_{2}c^{\prime\prime}_{1}-v^{\prime}_{1}c^{\prime\prime}_{3})^{2}+{v^{\prime}_{1}}^{2}{c^{\prime\prime}_{2}}^{2}\big]+\frac{1}{4}{c^{\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}_{s}}-\frac{1}{4}{v^{\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{CB}_{p}},$ (149)
$\mathcal{V}_{\mathcal{CB}_{sp}}-\mathcal{V}_{\mathcal{N}_{s}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime}}^{2}}\right)_{\mathcal{N}_{s}}\big[(v^{\prime}_{2}c^{\prime\prime\prime}_{1}-v^{\prime}_{1}c^{\prime\prime\prime}_{3})^{2}+{v^{\prime}_{1}}^{2}{c^{\prime\prime\prime}_{2}}^{2}\big]+\frac{1}{4}{c^{\prime\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}_{s}}+\frac{1}{2}c^{\prime\prime\prime}_{S}c^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{N}_{s}}$
$-\frac{1}{4}{v^{\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{CB}_{sp}},$ (150)
$\mathcal{V}_{\mathcal{CP}}-\mathcal{V}_{\mathcal{N}_{s}}=$ $\left(\frac{m_{A}^{2}}{4{v^{\prime}}^{2}}\right)_{\mathcal{N}_{s}}\big[(v^{\prime}_{2}\bar{v}_{1}-v^{\prime}_{1}\bar{v}_{2})^{2}+{v^{\prime}_{1}}^{2}\bar{v}_{3}^{2}\big]-\frac{1}{4}{v^{\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{CP}},$ (151)
$\mathcal{V}_{\mathcal{CP}_{s}}-\mathcal{V}_{\mathcal{N}_{s}}=$ $\left(\frac{m_{A}^{2}}{4{v^{\prime}}^{2}}\right)_{\mathcal{N}_{s}}\big[(v^{\prime}_{2}\bar{v}^{\prime}_{1}-v^{\prime}_{1}\bar{v}^{\prime}_{2})^{2}+{v^{\prime}_{1}}^{2}{\left.\bar{v}^{\prime}_{3}\right.}^{2}\big],$ (152)
$\mathcal{V}_{\mathcal{CP}_{p}}-\mathcal{V}_{\mathcal{N}_{s}}=$ $\left(\frac{m_{A}^{2}}{4{v^{\prime}}^{2}}\right)_{\mathcal{N}_{s}}\big[(v^{\prime}_{2}\bar{v}^{\prime\prime}_{1}-v^{\prime}_{1}\bar{v}^{\prime\prime}_{2})^{2}+{v^{\prime}_{1}}^{2}{\left.\bar{v}^{\prime\prime}_{3}\right.}^{2}\big]+\frac{1}{4}{\left.\bar{v}^{\prime\prime}_{P}\right.}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}_{s}}-\frac{1}{4}{v^{\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{CP}_{p}},$ (153)
$\mathcal{V}_{\mathcal{CP}_{sp}}-\mathcal{V}_{\mathcal{N}_{s}}=$ $\left(\frac{m_{A}^{2}}{4{v^{\prime}}^{2}}\right)_{\mathcal{N}_{s}}\big[(v^{\prime}_{2}\bar{v}^{\prime\prime\prime}_{1}-v^{\prime}_{1}\bar{v}^{\prime\prime\prime}_{2})^{2}+{v^{\prime}_{1}}^{2}{\left.\bar{v}^{\prime\prime\prime}_{3}\right.}^{2}\big]+\frac{1}{4}{\left.\bar{v}^{\prime\prime\prime}_{P}\right.}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}_{s}}+\frac{1}{2}\bar{v}^{\prime\prime\prime}_{S}\bar{v}^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{N}_{s}}$
$-\frac{1}{4}{v^{\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{CP}_{sp}}.$ (154)
From these expressions, we conclude that if the potential has a minimum of type $\mathcal{N}_{s}$, any stationary point of type $\mathcal{CB}_{s}$ or $\mathcal{CP}_{s}$, if it exists, lies above $\mathcal{N}_{s}$. This is not necessarily the case for the other stationary points of type $\mathcal{CB}$ and $\mathcal{CP}$.
A.3 $\mathcal{N}_{p}$-type vacuum
The expressions for comparing the depth of the $\mathcal{N}_{p}$-type potential against $\mathcal{CB}$ and $\mathcal{CP}$-type vacua are:
$\mathcal{V}_{\mathcal{CB}}-\mathcal{V}_{\mathcal{N}_{p}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime}}^{2}}\right)_{\mathcal{N}_{p}}\big[(v^{\prime\prime}_{2}c_{1}-v^{\prime\prime}_{1}c_{3})^{2}+{v^{\prime\prime}_{1}}^{2}c_{2}^{2}\big]-\frac{1}{4}{v^{\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{CB}},$ (155)
$\mathcal{V}_{\mathcal{CB}_{s}}-\mathcal{V}_{\mathcal{N}_{p}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime}}^{2}}\right)_{\mathcal{N}_{p}}\big[(v^{\prime\prime}_{2}c^{\prime}_{1}-v^{\prime\prime}_{1}c^{\prime}_{3})^{2}+{v^{\prime\prime}_{1}}^{2}{c^{\prime}_{2}}^{2}\big]+\frac{1}{4}{c^{\prime}_{S}}^{2}\big(m_{S}^{2}\big)_{\mathcal{N}_{p}}-\frac{1}{4}{v^{\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{CB}_{s}},$ (156)
$\mathcal{V}_{\mathcal{CB}_{p}}-\mathcal{V}_{\mathcal{N}_{p}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime}}^{2}}\right)_{\mathcal{N}_{p}}\big[(v^{\prime\prime}_{2}c^{\prime\prime}_{1}-v^{\prime\prime}_{1}c^{\prime\prime}_{3})^{2}+{v^{\prime\prime}_{1}}^{2}{c^{\prime\prime}_{2}}^{2}\big],$ (157)
$\mathcal{V}_{\mathcal{CB}_{sp}}-\mathcal{V}_{\mathcal{N}_{p}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime}}^{2}}\right)_{\mathcal{N}_{p}}\big[(v^{\prime\prime}_{2}c^{\prime\prime\prime}_{1}-v^{\prime\prime}_{1}c^{\prime\prime\prime}_{3})^{2}+{v^{\prime\prime}_{1}}^{2}{c^{\prime\prime\prime}_{2}}^{2}\big]+\frac{1}{4}{c^{\prime\prime\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}_{p}}+\frac{1}{2}c^{\prime\prime\prime}_{S}c^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{N}_{p}}$
$-\frac{1}{4}{v^{\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{CB}_{sp}},$ (158)
$\mathcal{V}_{\mathcal{CP}}-\mathcal{V}_{\mathcal{N}_{p}}=$ $\left(\frac{m_{A}^{2}}{4{v^{\prime\prime}}^{2}}\right)_{\mathcal{N}_{p}}\big[(v^{\prime\prime}_{2}\bar{v}_{1}-v^{\prime\prime}_{1}\bar{v}_{2})^{2}+{v^{\prime\prime}_{1}}^{2}\bar{v}_{3}^{2}\big]-\frac{1}{4}{v^{\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{CP}},$ (159)
$\mathcal{V}_{\mathcal{CP}_{s}}-\mathcal{V}_{\mathcal{N}_{p}}=$ $\left(\frac{m_{A}^{2}}{4{v^{\prime\prime}}^{2}}\right)_{\mathcal{N}_{p}}\big[(v^{\prime\prime}_{2}\bar{v}^{\prime}_{1}-v^{\prime\prime}_{1}\bar{v}^{\prime}_{2})^{2}+{v^{\prime\prime}_{1}}^{2}{\left.\bar{v}^{\prime}_{3}\right.}^{2}\big]+\frac{1}{4}{\left.\bar{v}^{\prime}_{S}\right.}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}_{p}}-\frac{1}{4}{v^{\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{CP}_{s}},$ (160)
$\mathcal{V}_{\mathcal{CP}_{p}}-\mathcal{V}_{\mathcal{N}_{p}}=$ $\left(\frac{m_{A}^{2}}{4{v^{\prime\prime}}^{2}}\right)_{\mathcal{N}_{p}}\big[(v^{\prime\prime}_{2}\bar{v}^{\prime\prime}_{1}-v^{\prime\prime}_{1}\bar{v}^{\prime\prime}_{2})^{2}+{v^{\prime\prime}_{1}}^{2}{\left.\bar{v}^{\prime\prime}_{3}\right.}^{2}\big],$ (161)
$\mathcal{V}_{\mathcal{CP}_{sp}}-\mathcal{V}_{\mathcal{N}_{p}}=$ $\left(\frac{m_{A}^{2}}{4{v^{\prime\prime}}^{2}}\right)_{\mathcal{N}_{p}}\big[(v^{\prime\prime}_{2}\bar{v}^{\prime\prime\prime}_{1}-v^{\prime\prime}_{1}\bar{v}^{\prime\prime\prime}_{2})^{2}+{v^{\prime\prime}_{1}}^{2}{\left.\bar{v}^{\prime\prime\prime}_{3}\right.}^{2}\big]+\frac{1}{4}{\left.\bar{v}^{\prime\prime\prime}_{S}\right.}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}_{p}}+\frac{1}{2}\bar{v}^{\prime\prime\prime}_{S}\bar{v}^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{N}_{p}}$
$-\frac{1}{4}{v^{\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{CP}_{sp}}.$ (162)
Analogously to the previous case, if the potential has a minimum of type $\mathcal{N}_{p}$, any stationary point of type $\mathcal{CB}_{p}$ or $\mathcal{CP}_{p}$, if it exists, lies above $\mathcal{N}_{p}$. This is not necessarily the case for the other stationary points of type $\mathcal{CB}$ and $\mathcal{CP}$.
A.4 $\mathcal{N}_{sp}$-type vacuum
The expressions for comparing the depth of the $\mathcal{N}_{sp}$-type potential against $\mathcal{CB}$ and $\mathcal{CP}$-type vacua are:
$\mathcal{V}_{\mathcal{CB}}-\mathcal{V}_{\mathcal{N}_{sp}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime\prime}}^{2}}\right)_{\mathcal{N}_{sp}}\big[(v^{\prime\prime\prime}_{2}c_{1}-v^{\prime\prime\prime}_{1}c_{3})^{2}+{v^{\prime\prime\prime}_{1}}^{2}c_{2}^{2}\big]-\frac{1}{4}{v^{\prime\prime\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{CB}}-\frac{1}{4}{v^{\prime\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{CB}}$
$-\frac{1}{2}v^{\prime\prime\prime}_{S}v^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{CB}},$ (163)
$\mathcal{V}_{\mathcal{CB}_{s}}-\mathcal{V}_{\mathcal{N}_{sp}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime\prime}}^{2}}\right)_{\mathcal{N}_{sp}}\big[(v^{\prime\prime\prime}_{2}c^{\prime}_{1}-v^{\prime\prime\prime}_{1}c^{\prime}_{3})^{2}+{v^{\prime\prime\prime}_{1}}^{2}{c^{\prime}_{2}}^{2}\big]+\frac{1}{4}{c^{\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}_{sp}}-\frac{1}{4}{v^{\prime\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{CB}_{s}}$
$-\frac{1}{2}v^{\prime\prime\prime}_{S}v^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{CB}_{s}},$ (164)
$\mathcal{V}_{\mathcal{CB}_{p}}-\mathcal{V}_{\mathcal{N}_{sp}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime\prime}}^{2}}\right)_{\mathcal{N}_{sp}}\big[(v^{\prime\prime\prime}_{2}c^{\prime\prime}_{1}-v^{\prime\prime\prime}_{1}c^{\prime\prime}_{3})^{2}+{v^{\prime\prime\prime}_{1}}^{2}{c^{\prime\prime}_{2}}^{2}\big]+\frac{1}{4}{c^{\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}_{sp}}-\frac{1}{4}{v^{\prime\prime\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{CB}_{p}}$
$-\frac{1}{2}v^{\prime\prime\prime}_{S}v^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{CB}_{p}},$ (165)
$\mathcal{V}_{\mathcal{CB}_{sp}}-\mathcal{V}_{\mathcal{N}_{sp}}=$ $\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime\prime}}^{2}}\right)_{\mathcal{N}_{sp}}\big[(v^{\prime\prime\prime}_{2}c^{\prime\prime\prime}_{1}-v^{\prime\prime\prime}_{1}c^{\prime\prime\prime}_{3})^{2}+{v^{\prime\prime\prime}_{1}}^{2}{c^{\prime\prime\prime}_{2}}^{2}\big]+\frac{1}{4}{c^{\prime\prime\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}_{sp}}+\frac{1}{4}{c^{\prime\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}_{sp}}$
$+\frac{1}{2}c^{\prime\prime\prime}_{S}c^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{N}_{sp}}-\frac{1}{4}{v^{\prime\prime\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{CB}_{sp}}-\frac{1}{4}{v^{\prime\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{CB}_{sp}}-\frac{1}{2}v^{\prime\prime\prime}_{S}v^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{CB}_{p}},$ (166)
$\mathcal{V}_{\mathcal{CP}}-\mathcal{V}_{\mathcal{N}_{sp}}=$ $\left(\frac{m_{A}^{2}}{4{v^{\prime\prime\prime}}^{2}}\right)_{\mathcal{N}_{sp}}\big[(v^{\prime\prime\prime}_{2}\bar{v}_{1}-v^{\prime\prime\prime}_{1}\bar{v}_{2})^{2}+{v^{\prime\prime\prime}_{1}}^{2}\bar{v}_{3}^{2}\big]-\frac{1}{4}{v^{\prime\prime\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{CP}}-\frac{1}{4}{v^{\prime\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{CP}}$
$-\frac{1}{2}v^{\prime\prime\prime}_{S}v^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{CP}},$ (167)
$\mathcal{V}_{\mathcal{CP}_{s}}-\mathcal{V}_{\mathcal{N}_{sp}}=$ $\left(\frac{m_{A}^{2}}{4{v^{\prime\prime\prime}}^{2}}\right)_{\mathcal{N}_{sp}}\big[(v^{\prime\prime\prime}_{2}\bar{v}^{\prime}_{1}-v^{\prime\prime\prime}_{1}\bar{v}^{\prime}_{2})^{2}+{v^{\prime\prime\prime}_{1}}^{2}{\left.\bar{v}^{\prime}_{3}\right.}^{2}\big]+\frac{1}{4}{\left.\bar{v}^{\prime}_{S}\right.}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}_{sp}}-\frac{1}{4}{v^{\prime\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{CP}_{s}}$
$-\frac{1}{2}v^{\prime\prime\prime}_{S}v^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{CP}_{s}},$ (168)
$\mathcal{V}_{\mathcal{CP}_{p}}-\mathcal{V}_{\mathcal{N}_{sp}}=$ $\left(\frac{m_{A}^{2}}{4{v^{\prime\prime\prime}}^{2}}\right)_{\mathcal{N}_{sp}}\big[(v^{\prime\prime\prime}_{2}\bar{v}^{\prime\prime}_{1}-v^{\prime\prime\prime}_{1}\bar{v}^{\prime\prime}_{2})^{2}+{v^{\prime\prime\prime}_{1}}^{2}{\left.\bar{v}^{\prime\prime}_{3}\right.}^{2}\big]+\frac{1}{4}{\left.\bar{v}^{\prime\prime}_{P}\right.}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}_{sp}}-\frac{1}{4}{v^{\prime\prime\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{CP}_{p}}$
$-\frac{1}{2}v^{\prime\prime\prime}_{S}v^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{CP}_{p}},$ (169)
$\mathcal{V}_{\mathcal{CP}_{sp}}-\mathcal{V}_{\mathcal{N}_{sp}}=$ $\left(\frac{m_{A}^{2}}{4{v^{\prime\prime\prime}}^{2}}\right)_{\mathcal{N}_{sp}}\big[(v^{\prime\prime\prime}_{2}\bar{v}^{\prime\prime\prime}_{1}-v^{\prime\prime\prime}_{1}\bar{v}^{\prime\prime\prime}_{2})^{2}+{v^{\prime\prime\prime}_{1}}^{2}{\left.\bar{v}^{\prime\prime\prime}_{3}\right.}^{2}\big]+\frac{1}{4}{\left.\bar{v}^{\prime\prime\prime}_{S}\right.}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}_{sp}}+\frac{1}{4}{\left.\bar{v}^{\prime\prime\prime}_{P}\right.}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}_{sp}}$
$+\frac{1}{2}\bar{v}^{\prime\prime\prime}_{S}\bar{v}^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{N}_{sp}}-\frac{1}{4}{v^{\prime\prime\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{CP}_{sp}}-\frac{1}{4}{v^{\prime\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{CP}_{sp}}-\frac{1}{2}v^{\prime\prime\prime}_{S}v^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{CP}_{p}}.$ (170)
There is no guarantee that the $\mathcal{N}_{sp}$ vacuum is deeper than any $\mathcal{CB}$ or $\mathcal{CP}$-type vacuum.
A.5 Coexisting Neutral minima
Another contingency that we must take is to verify the stability between $\mathcal{N},\mathcal{N}_{s},\mathcal{N}_{p}$ and $\mathcal{N}_{sp}$. The equations relating their potential depth are the following.
$\mathcal{V}_{\mathcal{N}_{s}}-\mathcal{V}_{\mathcal{N}}=$ $\left[\left(\frac{m_{H^{\pm}}^{2}}{4v^{2}}\right)_{\mathcal{N}}-\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime}}^{2}}\right)_{\mathcal{N}_{s}}\right](v_{2}^{\prime}v_{1}-v_{1}^{\prime}v_{2})^{2}+\frac{1}{4}{v^{\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}},$ (171)
$\mathcal{V}_{\mathcal{N}_{p}}-\mathcal{V}_{\mathcal{N}}=$ $\left[\left(\frac{m_{H^{\pm}}^{2}}{4v^{2}}\right)_{\mathcal{N}}-\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime}}^{2}}\right)_{\mathcal{N}_{p}}\right](v_{2}^{\prime\prime}v_{1}-v_{1}^{\prime\prime}v_{2})^{2}+\frac{1}{4}{v^{\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}},$ (172)
$\mathcal{V}_{\mathcal{N}_{sp}}-\mathcal{V}_{\mathcal{N}}=$ $\left[\left(\frac{m_{H^{\pm}}^{2}}{4v^{2}}\right)_{\mathcal{N}}-\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime\prime}}^{2}}\right)_{\mathcal{N}_{sp}}\right](v_{2}^{\prime\prime\prime}v_{1}-v_{1}^{\prime\prime\prime}v_{2})^{2}+\frac{1}{4}{v^{\prime\prime\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}}+\frac{1}{4}{v^{\prime\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}}$
$+\frac{1}{2}v^{\prime\prime\prime}_{S}v^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{N}},$ (173)
$\mathcal{V}_{\mathcal{N}_{p}}-\mathcal{V}_{\mathcal{N}_{s}}=$ $\left[\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime}}^{2}}\right)_{\mathcal{N}_{s}}-\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime}}^{2}}\right)_{\mathcal{N}_{p}}\right](v_{2}^{\prime\prime}v_{1}^{\prime}-v_{1}^{\prime\prime}v_{2}^{\prime})^{2}+\frac{1}{4}{v^{\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}_{s}}-\frac{1}{4}{v^{\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}_{p}},$ (174)
$\mathcal{V}_{\mathcal{N}_{sp}}-\mathcal{V}_{\mathcal{N}_{s}}=$ $\left[\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime}}^{2}}\right)_{\mathcal{N}_{s}}-\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime\prime}}^{2}}\right)_{\mathcal{N}_{sp}}\right](v_{2}^{\prime\prime\prime}v_{1}^{\prime}-v_{1}^{\prime\prime\prime}v_{2}^{\prime})^{2}+\frac{1}{4}{v^{\prime\prime\prime}_{P}}^{2}\big(m_{p}^{2}\big)_{\mathcal{N}_{s}}+\frac{1}{2}v^{\prime\prime\prime}_{S}v^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{N}_{s}}$
$-\frac{1}{4}{v^{\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}_{sp}},$ (175)
$\mathcal{V}_{\mathcal{N}_{sp}}-\mathcal{V}_{\mathcal{N}_{p}}=$ $\left[\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime}}^{2}}\right)_{\mathcal{N}_{p}}-\left(\frac{m_{H^{\pm}}^{2}}{4{v^{\prime\prime\prime}}^{2}}\right)_{\mathcal{N}_{sp}}\right](v_{2}^{\prime\prime\prime}v_{1}^{\prime\prime}-v_{1}^{\prime\prime\prime}v_{2}^{\prime\prime})^{2}+\frac{1}{4}{v^{\prime\prime\prime}_{S}}^{2}\big(m_{s}^{2}\big)_{\mathcal{N}_{p}}+\frac{1}{2}v^{\prime\prime\prime}_{S}v^{\prime\prime\prime}_{P}\big(m_{sp}^{2}\big)_{\mathcal{N}_{p}}$
$-\frac{1}{4}{v^{\prime\prime}_{P}}^{2}\big({m_{p}}^{2}\big)_{\mathcal{N}_{sp}}.$ (176)
From them, it is possible to infer that if the $\mathcal{N},\mathcal{N}_{s}$ and $\mathcal{N}_{p}$ minima coexist in the potential, then the global minimum conserves charge and CP, but if $\mathcal{N}_{sp}$ also exists, the same may not always be true.
The minima of the types $\mathcal{S}$ and $\mathcal{P}$ exist only if $m^{2}_{S}<0$ and $m^{2}_{P}<0$. Hence, the analysis of their stability must be done numerically. It is also not possible to analyze the stability of the $\mathcal{SP}$ minima analytically.